### v5

Abstract
For most problems pertaining to perfect matchings, one may restrict attention to matching covered graphs — that is, connected nontrivial graphs with the property that each edge belongs to some perfect matching. There is extensive literature on these graphs that are also known as $1$-extendable graphs (since each edge extends to a perfect matching) including an ear decomposition theorem due to Lovász and Plummer.
A cycle $C$ of a graph $G$ is conformal if $G-V(C)$ has a perfect matching; such cycles play an important role in the study of perfect matchings, especially when investigating the Pfaffian orientation problem. A matching covered graph $G$ is cycle-extendable if — for each even cycle $C$ — the cycle $C$ is conformal, or equivalently, each perfect matching of $C$ extends to a perfect matching of $G$, or equivalently, $C$ is the symmetric difference of two perfect matchings of $G$, or equivalently, $C$ extends to an ear decomposition of $G$. In the literature, these are also known as cycle-nice or as $1$-cycle resonant graphs.
Zhang, Wang, Yuan, Ng and Cheng, 2022, provided a characterization of claw-free cycle-extendable graphs. Guo and Zhang, 2004, and independently Zhang and Li, 2012, provided characterizations of bipartite planar cycle-extendable graphs. In this paper, we establish a characterization of all planar cycle-extendable graphs — in terms of $K_{2}$ and four infinite families.
keywords:
matchings, perfect matchings, matching covered graphs, cycle-extendability
1 Cycle-extendability
All graphs considered in this paper are loopless. However, we allow multiple/parallel edges. For graph-theoretic notation and terminology, we follow Bondy and Murty (2008), whereas for terminology pertaining to matching theory, we follow Lucchesi and Murty (2024).
A graph is said to be matchable if it has a perfect matching. Most problems pertaining to the study of perfect matchings may be reduced to matching covered graphs — that is, nontrivial connected graphs with the property that each edge lies in some perfect matching. There is extensive literature on matching covered graphs. In Lovász and Plummer (1986), they are referred to as $1$-extendable graphs, since each edge extends to a perfect matching. There is also literature on $k$-extendable graphs — matching covered graphs with the additional property that each matching of cardinality $k$ extends to a perfect matching; see Plummer (1991). In a similar spirit, a matching covered graph $G$ is cycle-extendable if, for each even cycle $C$, either perfect matching of $C$ extends to a perfect matching of $G$. This leads us to the following decision problem that is not known to be in NP.
Decision Problem 0.1
Given a matching covered graph $G$, decide whether $G$ is cycle-extendable.
Before stating our contributions and related prior work on the above problem, let us take a closer look at conformality and cycle-extendability. A cycle $C$ of a matchable graph $G$ is a conformal cycle if the graph $G-V(C)$ is matchable. Thus, a matching covered graph is cycle-extendable if and only if each even cycle is conformal. From this viewpoint, it is easy to see that Decision Problem 0.1 belongs to co-NP.
In Figure 1(a), the $6$-cycle denoted by the dashed line is not conformal, and using the symmetries of the Cube graph, the reader may verify that there are precisely four such $6$-cycles , whereas every other cycle is conformal. In particular, the Cube graph is not cycle-extendable. Figures 1(b) and 1(c) depict examples of bipartite and nonbipartite cycle-extendable graphs, respectively.
$v$
Figure 1: (a) the Cube graph is not cycle-extendable; (b) a bipartite cycle-extendable graph; (c) a nonbipartite cycle-extendable graph $W_{5}^{-}$
Observe that, in a matchable graph, a cycle is conformal if and only if it may be expressed as the symmetric difference of two perfect matchings. It is for this reason that they are also known as alternating cycles; see Carvalho and Little (2008). It is easily observed that a connected matchable graph, distinct from $K_{2}$, is matching covered if and only if each edge belongs to a conformal cycle. Little (1975) proved the stronger statement that, in a matching covered graph, any two edges belong to a common conformal cycle. There are several other terms used in the literature for conformal cycles : well-fitted cycles in McCuaig (2004); nice cycles in Lovász (1987); and central cycles in Robertson et al. (1999).
Zhang, Wang, Yuan, Ng and Cheng Zhang et al. (2022) provided a characterization of claw-free cycle-extendable (aka cycle-nice) matching covered graphs; their work implies that Decision Problem 0.1 belongs to NP as well as P for claw-free graphs.
Guo and Zhang (2004), and independently Zhang and Li (2012), provided characterizations of bipartite planar cycle-extendable (aka 1-cycle resonant) graphs. The second author discovered a couple of these characterizations independently, and they appear in their MSc thesis Pause (2022); they also made partial progress towards characterizing planar cycle-extendable graphs.
In this paper, we characterize all planar cycle-extendable graphs and our result implies that Decision Problem 0.1 belongs to NP as well as P for planar graphs. The results reported here comprise a proper subset of those that appear in the MS thesis Dalwadi (2025) of the first author.
Organization of this paper
In Section 1.1, we discuss the ear decomposition theory for matching covered graphs that provides an alternative definition of cycle-extendable graphs that motivates their investigation. In Section 1.2, we discuss series and parallel reductions that help us in restricting ourselves to irreducible graphs — that is, simple graphs whose vertices of degree two comprise a stable set.
In Section 2, we describe the tight cut decomposition theory for matching covered graphs and its applications towards characterizing cycle-extendable graphs. In particular, in Section 2.2, we use a result of  Carvalho and Little (2008) to deduce that every planar cycle-extendable irreducible graph, except $K_{2}$, either has a vertex of degree two or otherwise is a “brick” — a special class of $3$-connected nonbipartite matching covered graphs.
In Section 3, we first discuss a necessary condition ($K_{2,3}$-freeness) for a planar matching covered graph to be cycle-extendable, and then a necessary condition for a $3$-connected graph to be $K_{2,3}$-free. In Section 4, we use results of Section 3 and the well-known brick generation theorem of Norine and Thomas (2007) to characterize planar cycle-extendable bricks.
In Section 5, we describe four infinite families of nonbipartite planar cycle-extendable irreducible graphs using a special class of bipartite cycle-extendable graphs called “half biwheels”. Finally, in Section 5.7, we prove that every planar cycle-extendable irreducible graph is either $K_{2}$ or is a member of one of these four families; this proves a conjecture of the third author, and places Decision Problem 0.1 in NP as well as in P for planar graphs.
1.1 Ear decompositions of matching covered graphs
Observe that matching covered graphs, except $K_{2}$, are $2$-connected. The ear decomposition theory of $2$-connected graphs, due to Whitney (1933), is well-known. We now discuss a refinement of this theory, due to Hetyei, that is applicable to the subclass of bipartite matching covered graphs; see Lovász and Plummer (1986).
Given a graph $G$ and a (proper) subgraph $H$, an ear (aka a single ear) of $H$ in $G$ is an odd path $P$ whose ends are in $V(H)$ but is otherwise disjoint with $H$. For a bipartite graph $G$, a sequence of subgraphs $(G_{0},G_{1},\dots,G_{r})$ is called a bipartite ear decomposition of $G$ if (i) $G_{0}$ is an (even) cycle, (ii) $G_{r}=G$, and (iii) $G_{i+1}=G_{i}\cup P_{i}$ where $P_{i}$ is an ear of $G_{i}$ (in $G$) for each $i\in\{0,1,\ldots,r-1\}$. Figure 2 shows a bipartite ear decomposition of a bipartite matching covered graph where each ear is denoted by thick lines.
Figure 2: a bipartite ear decomposition of a bipartite matching covered graph
A subgraph $H$ of a graph $G$ is conformal if the graph $G-V(H)$ is matchable. Given any bipartite ear decomposition of a bipartite graph, it is easily observed that each subgraph in the sequence is conformal. Furthermore, the following theorem due to Hetyei, implies that each subgraph is also matching covered and establishes the converse as well.
Theorem 0.2
[Bipartite Ear Decomposition Theorem]
A bipartite graph $G$, distinct from $K_{2}$, is matching covered if and only if each of its conformal cycles extends to a bipartite ear decomposition of $G$.
For the Cube graph, depicted in Figure 1(a), the reader may observe that each conformal cycle extends to a bipartite ear decomposition; furthermore, the non-conformal $6$-cycle (shown using dashed lines) does not extend to a bipartite ear decomposition.
Theorem 0.2 implies that any bipartite matching covered graph (distinct from $K_{2}$) can be constructed in a straightforward manner, from any conformal cycle, by adding a single ear at a time so that each subgraph is also matching covered graph. Unfortunately, in the case of nonbipartite matching covered graphs, one can not restrict to the addition of single ears. For instance, in the case of $K_{4}$, we must start from $C_{4}$; now observe that we must add the remaining two edges simultaneously in order to get a bigger matching covered subgraph (that is, $K_{4}$ itself). Lovász and Plummer (1986) proved the surprising result that every matching covered graph may be constructed, from any conformal cycle, by adding either a single ear or a “double ear” at a time so that each subgraph is also matching covered. We state this more formally below.
Given a graph $G$ and (proper) subgraph $H$, a double ear of $H$ in $G$ is a pair of vertex-disjoint single ears of $H$ (in $G$). For a matching covered graph $G$, a sequence of matching covered subgraphs $(G_{0},G_{1},\dots,G_{r})$ is called an ear decomposition of $G$ if (i) $G_{0}$ is an even cycle, (ii) $G_{r}=G$, and (iii) $G_{i+1}=G_{i}\cup R_{i}$ where $R_{i}$ is either a single or a double ear of $G_{i}$ (in $G$) for each $i\in\{0,1,\ldots,r-1\}$. As in the case of bipartite ear decompositions, it is easy to see that each subgraph in an ear decomposition is also conformal. We are now ready to state the aforementioned theorem of Lovász and Plummer.
Theorem 0.3
[Ear Decomposition Theorem]
Each conformal cycle of a matching covered graph $G$ extends to an ear decomposition of $G$.
Figure 3 shows an ear decomposition of the nonbipartite graph $R_{8}^{-}$ that will play an important role in our work, and is obtained from the bicorn $R_{8}$, shown in Figure 6(a), by deleting an edge; each ear addition is denoted by a thick line.
Figure 3: an ear decomposition of $R_{8}^{-}$
The statement of Theorem 0.3 begs the question: for which matching covered graphs, is it possible to extend each even cycle to an ear decomposition? Clearly, this is possible only for those matching covered graphs that are cycle-extendable. This provides an alternative motivation for characterizing cycle-extendable graphs.
Proposition 0.4
[An Alternative Viewpoint Of Cycle-Extendability]
A matching covered graph $G$ is cycle-extendable if and only if each of its even cycles extends to an ear decomposition of $G$. $\qed$
We conclude this section by describing a special class of nonbipartite matching covered graphs that play a crucial role in the theory of matching covered graphs, as well as in our work. Before doing so, we remark that, in an ear decomposition of a matching covered graph, a double ear is added only when neither of the constituent single ears can be added in order to obtain a matching covered graph.
Let $G$ be a nonbipartite matching covered graph and $(G_{0},G_{1},G_{2},\ldots,G_{r}=G)$ be an ear decomposition. Note that $G_{0}$ is an even cycle. Let $G_{k}$ denote the first nonbipartite subgraph in this sequence. It follows from our above remark that $G_{k}$ is also the first subgraph in the sequence that is obtained by adding a double ear, say $R_{k-1}$, to the previous (bipartite) subgraph $G_{k-1}$. We say that $G_{k}$ is a near-bipartite graph. Below, we provide an alternative definition that is independent of the ear decomposition.
For a matching covered graph $G$, an odd path $P$ is a removable single ear of $G$ if each internal vertex of $P$ (if any) has degree two in $G$ and the graph $G-P$ is also matching covered. (Here, by $G-P$, we mean the graph obtained from $G$ by deleting all of the edges and internal vertices of $P$.) Likewise, a pair of vertex-disjoint odd paths $R:=\{P_{1},P_{2}\}$ is a removable double ear of $G$ if each internal vertex of $P_{1}$ as well as $P_{2}$ has degree two in $G$ and the graph $G-P_{1}-P_{2}$ is matching covered. A near-bipartite graph is a matching covered graph, say $G$, that has a removable double ear $R:=\{P_{1},P_{2}\}$ such that $G-R:=G-P_{1}-P_{2}$ is bipartite. In this sense, near-bipartite graphs comprise a subset of nonbipartite matching covered graphs whose members are closest to being bipartite.
In the next section, we discuss a reduction of the Decision Problem 0.1 to a more restricted class of matching covered graphs.
1.2 Irreducible graphs
Given a graph $G$ and a parallel edge $e$, we say that $G-e$ is obtained from $G$ by an application of parallel reduction. Observe that $G$ is matching covered if and only if $G-e$ is matching covered; furthermore, $G$ is cycle-extendable if and only if $G-e$ is cycle-extendable.
Let $G$ be a graph that has a path $P:=wxyz$ of length three, each of whose internal vertices, $x$ and $y$, is of degree two in $G$. Let $J$ denote the graph obtained from $G$ by replacing the path $P$ with a single edge $e$ joining $w$ and $z$. That is, $J:=G-P+e$. We say that $J$ is obtained from $G$ by an application of series reduction. The reader may observe that $G$ is matching covered if and only if $J$ is matching covered except when $J$ is $K_{2}$; furthermore, $G$ is cycle-extendable if and only if $J$ is cycle-extendable.
A graph $G$ is irreducible if it is simple and its degree two vertices comprise a stable set. The above observations prove the following.
Proposition 0.5
Let $H$ be an irreducible (matching covered) graph that is obtained from a matching covered graph $G$ by repeated applications of series and parallel reductions. Then $G$ is cycle-extendable if and only if $H$ is cycle-extendable. $\qed$
Thus, in order to settle the complexity status of Decision Problem 0.1, it suffices to focus on the following decision problem.
Decision Problem 0.6
Given an irreducible matching covered graph $G$, decide whether $G$ is cycle-extendable or not.
In the next section, we discuss a necessary matching-theoretic condition for a graph to be cycle-extendable. In order to do so, we shall require some concepts from the theory of matching covered graphs.
2 Applications of the tight cut decomposition theory
For a nonempty proper subset $X$ of the vertices of a graph $G$, we denote by $\partial(X)$ the cut associated with $X$, that is, the set of all edges that have one end in $X$ and the other end in $\overline{X}:=V(G)-X$. We refer to $X$ and $\overline{X}$ as the shores of $\partial(X)$. For a vertex $v$, we simplify the notation ($\partial(\{v\})$ to $\partial(v)$, and we refer to such a cut as trivial.
For a cut $\partial(X)$ of a graph $G$, we denote by $G/(X\rightarrow x)$, or simply by $G/X$, the graph obtained from $G$ by shrinking the shore $X$ to a single vertex $x$ called the contraction vertex. The graph $G/\overline{X}$ is defined analogously. The graphs $G/\overline{X}$ and $G/X$ are called the $\partial(X)$-contractions of $G$. Observe that each edge in either $\partial(X)$-contraction corresponds to an edge in $G$. Conversely, each edge of $G$ that is not in $\partial(X)$ corresponds to an edge in precisely one of the two $\partial(X)$-contractions, whereas each edge of $G$ that is in $\partial(X)$ corresponds to an edge in both $\partial(X)$-contractions; it is customary and convenient to use the same label for all these edges. For the graph shown in Figure 4, the blue line indicates a cut; one of the contractions is $K_{4}$, whereas the other contraction is obtained from $K_{3,3}$ by replacing any vertex by a triangle.
Figure 4: a matching covered graph and its nontrivial tight cuts
2.1 Tight cut decomposition, bricks and braces
Let $G$ be a matching covered graph. A cut $\partial(X)$ is a tight cut if $|M\cap\partial(X)|=1$ for every perfect matching $M$. In Figure 4, the blue lines indicate nontrivial tight cuts. It is easy to see that if $\partial(X)$ is a nontrivial tight cut then each $\partial(X)$-contraction is a matching covered graph that has fewer vertices than $G$. If either of the $\partial(X)$-contractions has a nontrivial tight cut, then that graph can be further decomposed into smaller matching covered graphs. We may repeat this procedure until we obtain a list of matching covered graphs — each of which is free of nontrivial tight cuts. This process is known as the tight cut decomposition procedure.
A matching covered graph free of nontrivial tight cuts is called a brace if it is bipartite; otherwise, it is called a brick. Thus, an application of the tight cut decomposition procedure to a matching covered graph results in a list of bricks and braces. For the graph shown in Figure 4, an application of the tight cut decomposition procedure yields two copies of $K_{4}$ and one copy of $K_{3,3}$. We now make a simple observation pertaining to bricks and braces.
Let $S:=\{u,v\}$ denote a $2$-vertex-cut in a matching covered graph $G$ on six or more vertices. Let $J$ denote a component of $G-S$. If $J$ is a nontrivial odd component then $\partial(V(J))$ is a nontrivial tight cut; whereas, if $J$ is an even component then $\partial(V(J)\cup\{u\})$ is a nontrivial tight cut. This immediately proves the following well-known fact.
Proposition 0.7
Every simple brick/brace, except $K_{2}$ and $C_{4}$, is $3$-connected.$\qed$
A matching covered graph may admit several applications of the tight cut decomposition procedure. However, Lovász (1987) proved the following remarkable result.
Theorem 0.8
[Unique Tight Cut Decomposition Theorem]
Any two applications of the tight cut decomposition procedure to a matching covered graph yield the same list of bricks and braces (up to multiplicities of edges).
The following characterization of planar cycle-extendable braces is an immediate consequence of Proposition 0.7 and a result of (Klavžar and Salem, 2012, Corollary 2.3).
Corollary 0.9
The only simple planar cycle-extendable braces are $K_{2}$ and $C_{4}$. $\qed$
We now switch our attention to proving the following lemma that relates cycle-extendability with tight cuts.
Lemma 0.10
Let $G$ be a matching covered graph and let $\partial(X)$ denote a nontrivial tight cut. If $G$ is cycle-extendable then both $\partial(X)$-contractions of $G$ are also cycle-extendable.
Proof 2.1.
We let $G_{1}:=G/(X\rightarrow x)$ and $G_{2}:=G/(\overline{X}\rightarrow\overline{x})$. Assume that $G$ is cycle-extendable. It suffices to prove that $G_{1}$ is cycle-extendable. To this end, let $Q_{1}$ denote an even cycle of $G_{1}$.
If $x\notin V(Q_{1})$ then $Q_{1}$ is a cycle of $G$. We let $M$ denote a perfect matching of $G-V(Q_{1})$. Observe that, since $M$ extends to a perfect matching of $G$, and since $\partial(X)$ is tight, $|M\cap\partial(X)|=1$. Consequently, $M_{1}:=M\cap E(G_{1})$ is a perfect matching of $G_{1}-V(Q_{1})$.
Now suppose that $x\in V(Q_{1})$, and let $e$ and $f$ denote the two edges of $Q_{1}$ incident with $x$. Let $Q_{2}$ denote an even cycle of $G_{2}$ containing $e$ and $f$. (To see why such a cycle exists, consider the symmetric difference of two perfect matchings of $G_{2}$: one containing $e$ and another containing $f$.) Observe that $Q:=Q_{1}\cup Q_{2}$ is an even cycle in $G$, and let $M$ denote a perfect matching of $G-V(Q)$. Note that $M\cap\partial(X)=\emptyset$ and that $M_{1}:=M\cap E(G_{1})$ is a perfect matching of $G_{1}-V(Q_{1})$.
In both cases, we have shown that $Q_{1}$ is a conformal cycle of $G_{1}$, whence $G_{1}$ is cycle-extendable. This completes the proof of Lemma 0.10.
It is worth noting that the converse of the above lemma does not hold in general. For instance, the graph shown in Figure 4 is not cycle-extendable since the cycle shown using a thick line is not conformal; however, its bricks and braces are cycle-extendable. Lemma 0.10 immediately yields the following.
Corollary 11.
Each brick and brace of a cycle-extendable graph is also cycle-extendable.∎
Lovázs’s theorem (0.8) leads to certain graph invariants that play a crucial role in the theory of matching covered graphs. In the following two sections, we discuss a couple of these invariants and related concepts that are relevant to our work.
2.2 Near-bricks versus near-bipartite graphs
It follows from the Unique Tight Cut Decomposition Theorem (0.8) that the number of bricks of a matching covered graph $G$ (obtained by any application of the tight cut decomposition procedure) is an invariant; we denote this by $b(G)$. For instance, $b(G)=2$ for the graph $G$ shown in Figure 4.
It is worth noting that $b(G)=0$ if and only if $G$ is bipartite. We say that $G$ is a near-brick if $b(G)=1$. In particular, every brick is a near-brick. The following is an easy consequence of Theorem 0.8 that we will find useful in the next section.
Corollary 12.
For a near-brick, given any nontrivial tight cut $\partial(X)$, one of the $\partial(X)$-contractions is bipartite and the other one is a near-brick. ∎
We now observe an easy refinement of the above corollary. Let $G$ be a near-brick that is not a brick and let $\partial(X)$ denote a nontrivial tight cut. By Corollary 12, one of the $\partial(X)$-contractions, say $G_{1}:=G/\overline{X}$, is bipartite. Let $Y$ be a minimal (not necessarily proper) subset of $X$ such that $|Y|\geq 3$ and $D:=\partial_{G_{1}}(Y)$ is a tight cut of $G_{1}$. By choice of $Y$, the graph $G_{1}/\overline{Y}$ is a brace, where $\overline{Y}:=V(G_{1})-Y$. Observe that $D$ is a nontrivial tight cut of $G$ and that one of the $D$-contractions of $G$ is isomorphic to $G_{1}/\overline{Y}$. This proves the following well-known fact.
Corollary 13.
In any near-brick that is not a brick, there exists a nontrivial tight cut $\partial(X)$ such that one of the $\partial(X)$-contractions is a brace and the other one is a near-brick.∎
Recall from Section 1.1 that a near-bipartite graph is a matching covered graph, say $G$, that has a removable double ear $R:=\{P_{1},P_{2}\}$ such that $G-R:=G-P_{1}-P_{2}$ is bipartite. The following result of Carvalho et al. (2002) implies that every near-bipartite graph is a near-brick.
Theorem 14.
Let $G$ be a matching covered graph, and let $R$ be a removable double ear. Then $b(G-R)=b(G)-1$.
The following result is an immediate consequence of Theorem 14, and it implies that every near-brick that has a removable double ear is near-bipartite. For an example, see the graph $R_{8}^{-}$ shown in Figure 3.
Corollary 15.
Let $G$ be a near-brick, and let $R$ be a removable double ear. Then $G-R$ is bipartite and matching covered. ∎
We remark that a near-brick need not be near-bipartite. For instance, wheels (defined in Section 4.1), of order six or more, are bricks that are not near-bipartite. This is easily seen using the next proposition — that follows from the facts: (i) the number of odd faces of a plane graph $G$, denoted by $\sf{f_{odd}}(G)$, is even, and (ii) deleting any edge may reduce $\sf{f_{odd}(G)}$ by at most two.
Proposition 16.
Every near-bipartite plane graph $G$ satisfies $\sf{f_{odd}}(G)\in\{2,4\}$. ∎
Lastly, we use $p(G)$ to denote the number of Petersen bricks of a matching covered graph $G$ (obtained by any application of the tight cut decomposition procedure), where Petersen brick refers to the Petersen graph up to multiple edges. In the next section, we apply a result of Carvalho and Little (2008) to deduce that every cycle-extendable graph is either bipartite or otherwise a near-brick whose unique brick is not a Petersen brick.
2.3 Cycle-extendable implies bipartite or near-brick
The aforementioned result of Carvalho and Little pertains to the cycle space $\mathcal{C}(G)$ and its various subspaces that are well-studied vector spaces; see Bondy and Murty (2008). We let $\mathcal{C}^{e}(G)$ denote the even space of a graph $G$ — that is, the subspace of the cycle space $\mathcal{C}(G)$ spanned by the even cycles. Likewise, $\mathcal{A}(G)$ denotes the alternating space — that is, the subspace of $\mathcal{C}^{e}(G)$ spanned by the conformal cycles.
For $2$-connected graphs, it is well-known that the $\mathcal{C}^{e}(G)$ is a proper subspace of $\mathcal{C}(G)$ if and only if $G$ is nonbipartite. In the same spirit, the following result of Carvalho and Little (2008) characterizes the matching covered graphs for which the alternating space $\mathcal{A}(G)$ is a proper subspace of the even space $\mathcal{C}^{e}(G)$ — in terms of the invariants $b$ and $p$.
Theorem 17.
For a matching covered graph $G$, the following are equivalent:
  1. (i)
    $\mathcal{A}(G)$ is a proper subspace of $\mathcal{C}^{e}(G)$,
  2. (ii)
    there exists an even cycle $C\in\mathcal{C}^{e}(G)-\mathcal{A}(G)$, and
  3. (iii)
    $b(G)+p(G)>1$.
Observe that, for a matching covered graph $G$, the inequality $b(G)+p(G)>1$ holds if and only if $G$ is nonbipartite, and either (i) $G$ is not a near-brick or otherwise (ii) $G$ is a near-brick whose unique brick is the Petersen brick. The following is an immediate consequence of the above theorem that we alluded to earlier.
Corollary 18.
[Cycle-Extendable implies Bipartite or Near-Brick]
Every cycle-extendable graph is either bipartite or otherwise a near-brick whose unique brick is not the Petersen brick.
Proof 2.2.
We shall prove the contrapositive. Let $G$ be a matching covered graph such that $b(G)+p(G)>1$. By Theorem 17, there exists an even cycle $C\in\mathcal{C}^{e}(G)-\mathcal{A}(G)$. Clearly, $C$ is not a conformal cycle; whence $G$ is not cycle-extendable.
We now deduce another consequence of the above corollary and some of the earlier results; in the case of bipartite graphs, we shall find a result of Chartrand, Kaugars and Lick useful.
Corollary 19.
[Minimum Degree Three or more implies Brick]
Every simple planar cycle-extendable graph, with minimum degree three or more, is a brick.
Proof 2.3.
Let $G$ be a simple planar cycle-extendable graph with $\delta(G)\geq 3$. First, suppose that $G$ is bipartite. Consider a planar embedding of $G$. Since $G$ is 2-connected and $\delta(G)\geq 3$, it follows from a result of Chartrand et al. (1972) that there exists a vertex $v$ so that $G-v$ is a 2-connected (plane) graph. We invoke Whitney’s well-known result Bondy and Murty (2008), and we let $C$ denote the facial (even) cycle that bounds the face containing the point corresponding to the deleted vertex $v$. By planarity, all neighbors of $v$ lie on $C$. Thus, $C$ is not conformal and $G$ is not cycle-extendable; contradiction.
Hence, $G$ is nonbipartite. By Corollary 18, $G$ is a near-brick. Suppose to the contrary that $G$ is not a brick. Using Lemma 13, there exists a nontrivial tight cut $\partial(X)$ so that one of the $\partial(X)$-contractions, say $G_{1}:=G/\overline{X}$, is a brace. We invoke Corollaries 11 and 0.9 to infer that $G_{1}$ is isomorphic to $C_{4}$ (up to multiple edges). Since $G$ is simple, the multiple edges of $G_{1}$ (if any) are incident with the contraction vertex. Let $v$ denote the unique vertex of $G_{1}$ that is not incident with any edge in $\partial(X)$; see Figure 5. Observe that $d_{G_{1}}(v)=2$ and that $d_{G}(v)=2$. This contradicts our hypothesis that $\delta(G)\geq 3$.
$v$$\partial(X)$$\overline{X}$
Figure 5: illustration for the proof of Theorem 19
Inspired by the proof of the above corollary in the bipartite case, we introduce the following definition that we shall find useful later. For a vertex $v$ of a graph $G$, a cycle $C$ (in $G-v$) is said to be $v$-isolating if $v$ is an isolated vertex in the graph $G-V(C)$.
In order to characterize planar cycle-extendable graphs, it follows from Proposition 0.5 and Corollary 19 that it suffices to characterize (i) planar cycle-extendable bricks, and (ii) planar cycle-extendable irreducible graphs that have a vertex of degree two. In the case of planar bricks, we shall solve a more general problem that we discuss in the next section.
3 $K_{2,3}$-freeness
To bisubdivide an edge means to subdivide it by inserting an even number of subdivision vertices. A graph $H$ is a bisubdivision of a graph $G$ if $H$ may be obtained from $G$ by selecting any (possibly empty) subset $F\subseteq E(G)$ and bisubdividing each edge in $F$, or equivalently, if $G$ may be obtained from $H$ by a sequence of series reductions.
Given a fixed graph $J$, a graph $G$ is said to be $J$-free if $G$ does not contain any subgraph $H$ that is a bisubdivision of $J$; otherwise, we say that $G$ is $J$-based. The bicorn $R_{8}$, shown in Figure 6(a), is $K_{2,3}$-based; the thick lines indicate a bisubdivision of $K_{2,3}$. The pentagonal prism minus a specific edge, shown in Figure 6(b), is $K_{2,3}$-free; the reader may verify this using Lemma 20 and Proposition 26.
Figure 6: (a) the bicorn $R_{8}$; (b) the pentagonal prism minus one edge
3.1 Planarity: cycle-extendable implies $K_{2,3}$-free
We now discuss a necessary condition for a planar graph to be cycle-extendable.
Lemma 20.
[Cycle-Extendability implies $K_{2,3}$-freeness in Planar Graphs]
If a planar matching covered graph is cycle-extendable then it is $K_{2,3}$-free.
Proof 3.1.
We prove the contrapositive. Let $G$ denote any planar embedding of a planar matching covered graph that is $K_{2,3}$-based, and let $H$ denote a subgraph that is a bisubdivision of $K_{2,3}$. Observe that $H$ comprises two cubic vertices, say $u$ and $v$, and three internally-disjoint even $uv$-paths, say $P_{1},P_{2}$ and $P_{3}$. See Figure 7. Any two of these paths, say $P_{i}$ and $P_{j}$ (where $i<j$), comprise a facial cycle of $H$, say $Q_{ij}$. We will argue that at least one of these three cycles is not conformal.
[…]
Figure 7: an illustration for the proof of Lemma 20
As per the Jordan Curve Theorem, each cycle $Q_{ij}$ partitions the rest of the plane into two regions: (i) the interior denoted by ${\sf int}(Q_{ij})$ and (ii) the exterior denoted by ${\sf ext}(Q_{ij})$. We adjust notation so that ${\sf int}(Q_{ij})$ refers to the region that does not meet the path $P_{k}$ (where $k$ is distinct from $i$ and $j$) as shown in Figure 7.
Now, assume without loss of generality that each of $Q_{12}$ and $Q_{23}$ is conformal. Since $Q_{12}$ is conformal, ${\sf int}(Q_{12})$ contains an even number of vertices. Likewise, ${\sf int}(Q_{23})$ contains an even number of vertices. Observe that ${\sf ext}(Q_{13})$ contains precisely all of the vertices that lie in ${\sf int}(Q_{12})\cup{\sf int}(Q_{23})$, plus the internal vertices of the path $P_{2}$. In particular, ${\sf ext}(Q_{13})$ contains an odd number of vertices; ergo, $Q_{13}$ is not conformal. Thus $G$ is not cycle-extendable. This completes the proof of Lemma 20.
The above lemma was proved by the second author and appeares in his MSc thesis Pause (2022). However, we learnt later this was already observed by Zhang and Li (2012).
It is worth noting that the converse of the above lemma does not hold in general. Figure 6(b) shows an example of a (nonbipartite) planar matching covered graph that is $K_{2,3}$-free but not cycle-extendable; the even cycle depicted by a thick line is not conformal. However, the converse indeed holds in the case of bipartite graphs; this was proved by Zhang and Li (2012), and independently by the second author Pause (2022). This inspires the following decision problem.
Decision Problem 21
Given a matching covered graph $G$, decide whether $G$ is $K_{2,3}$-free.
Thus, in the context of planar nonbipartite matching covered graphs, cycle-extendable graphs comprise a proper subset of $K_{2,3}$-free graphs. We are currently working on characterizing $K_{2,3}$-free matching covered graphs, and have managed to solve the problem in the case of $3$-connected graphs.
In light of Lemma 20, in order to conclude that a planar graph is not cycle-extendable, it suffices to show the existence of a bisubdivision of $K_{2,3}$. In the following section, we discuss a sufficient condition for the existence of a bisubdivision of $K_{2,3}$ that is applicable to $3$-connected nonbipartite graphs. Since bricks are $3$-connected (see Proposition 0.7) and nonbipartite, this will help us in characterizing planar cycle-extendable bricks. In fact, we will obtain a characterization of planar $K_{2,3}$-free bricks.
3.2 $3$-connectedness: mixed bicycle implies $K_{2,3}$-based
A pair of vertex-disjoint simple cycles, say $(Q_{o},Q_{e})$, is said to be a mixed bicycle if $Q_{o}$ is odd and $Q_{e}$ is even. (Here, simple just means that the length of $Q_{e}$ is four or more.) The following is our promised sufficient condition for $3$-connected (nonbipartite) graphs.
Lemma 22.
[Mixed Bicycle implies $K_{2,3}$-based in $3$-connected Graphs]
If a $3$-connected graph has a mixed bicycle then it is $K_{2,3}$-based.
Proof 3.2.
Let $(Q_{o},Q_{e})$ denote a mixed bicycle in a $3$-connected graph $G$. By Menger’s Theorem, there exist three vertex-disjoint paths — say $P_{i}$ (where $i\in\{1,2,3\}$) — each of which has one end, say $u_{i}$, in $Q_{o}$ and the other end, say $v_{i}$, in $Q_{e}$. Let $Q_{1},Q_{2}$ and $Q_{3}$ denote the three edge-disjoint paths of $Q_{e}$ that have both ends in $\{v_{1},v_{2},v_{3}\}$. In particular, $Q_{e}=Q_{1}\cup Q_{2}\cup Q_{3}$.
Since $Q_{e}$ is even, at least one of $Q_{1},Q_{2}$ and $Q_{3}$ is even. Adjust notation so that $Q_{1}$ is even, and let $v_{1}$ and $v_{2}$ denote the ends of $Q_{1}$. Note that $Q_{1}$ and $Q_{2}\cup Q_{3}$ are edge-disjoint even $v_{1}v_{2}$-paths. It remains to find one more even $v_{1}v_{2}$-path that is edge-disjoint with $Q_{1}\cup(Q_{2}\cup Q_{3})$. Such a path is easily obtained by combining $P_{1}\cup P_{2}$ with the $u_{1}u_{2}$-path of $Q_{o}$ of appropriate parity. This completes the proof.
We remark that the converse of the above lemma does not hold in general. Clearly, $3$-connected bipartite matching covered graphs are counterexamples; however, one can also construct nonbipartite counterexamples such as certain subgraphs of $K_{5}$ and $K_{6}$. We conclude this section with the following consequence of Lemmas 22 and 20.
Corollary 23.
If a $3$-connected planar matching covered graph has a mixed bicycle then it is not cycle-extendable. ∎
4 Planar $K_{2,3}$-free bricks
Our characterization of planar $K_{2,3}$-free bricks relies on two main ingredients, one of which is Lemma 22. The other one is the brick generation theorem due to Norine and Thomas (2007), which states that all simple bricks may be constructed from five infinite families and the Petersen graph by means of four operations. We refer to these families as Norine-Thomas families; furthermore, we refer to their members, as well as to the Petersen graph, as Norine-Thomas bricks.
In the next section, we discuss the planar Norine-Thomas families, and we use Lemma 22 to classify the $K_{2,3}$-free ones. This classification will serve as the base case in our inductive proof of the characterization of planar $K_{2,3}$-free bricks.
4.1 Planar Norine-Thomas bricks
We begin this section by describing two of the five Norine-Thomas families — odd wheels and odd prisms — whose members happen to be planar as well as cycle-extendable.
A graph obtained from an odd cycle $Q:=u_{0}u_{1}\ldots u_{2k}$, by adding a new vertex $h$ and edges $hu_{i}$ for each $i\in\{0,\ldots,2k\}$, is called an odd wheel, or simply a wheel. The vertex $h$ is called its hub and the cycle $Q$ is called its rim. Let $G$ be a wheel and let $C$ denote an even cycle. Clearly, $h\in V(C)$ and $G-V(C)$ is an odd path, whence $C$ is conformal. This proves the following.
Proposition 24.
Wheels are cycle-extendable.∎
A graph obtained from two vertex-disjoint copies of an odd cycle, say $Q:=w_{0}w_{1}\ldots w_{2k}$ and $Q^{{}^{\prime}}:=z_{0}z_{1}\ldots z_{2k}$, by adding edges $w_{i}z_{i}$ for each $i\in\{0,1,\ldots,2k\}$, is called an odd prism, or simply a prism. Observe that prisms are cubic — that is, the degree of each vertex is precisely three.
In order to prove that prisms are cycle-extendable, we will first observe that they are near-bipartite and we shall locate all of their removable double ears; the reader may recall definitions from Section 2.2. Before that, we state a simplification of the notion of removable double ears that is applicable to irreducible graphs.
For a matching covered graph $G$, an edge $e$ is removable if $G-e$ is matching covered, and a pair of distinct edges $R:=\{\alpha,\beta\}$ is a removable doubleton if neither $\alpha$ nor $\beta$ is removable but $G-R$ is matching covered. For instance, the graph $R_{8}^{-}$, shown in Figure 3, has a removable doubleton depicted by thick lines. Note that, if $G$ is irreducible, each removable double ear is a removable doubleton and vice-versa. Furthermore, if $G$ is near-bipartite then, by Corollary 15, the graph $G-R$ is bipartite for each removable doubleton $R:=\{\alpha,\beta\}$; also, if $A$ and $B$ denote the color classes of $G-R$ then (up to relabeling) the edge $\alpha$ has both ends in $A$ whereas $\beta$ has both ends in $B$. Using these observations, the following is easily proved; we shall it find very useful in future sections, and we shall invoke it implicitly.
Lemma 25.
For a near-bipartite graph $G$, if $R$ is any removable doubleton and $C$ is any even cycle, then $|C\cap R|\in\{0,2\}$. ∎
Let $G$ be a prism with vertex-disjoint odd cycles $Q:=w_{0}w_{1}\ldots w_{2k}$ and $Q^{{}^{\prime}}:=z_{0}z_{1}\ldots z_{2k}$ as in the definition stated earlier. Observe that $\{w_{i}w_{i+1},z_{i}z_{i+1}\}$ is a removable doubleton of $G$ for each $i\in\{0,1,\ldots,2k\}$, where arithmetic is modulo $2k$. In particular, each vertex is incident with two distinct removable doubletons. Using this observation and Lemma 25, and the fact that $G$ is cubic, we infer that an even cycle $C$ of $G$ contains $w_{i}$ if and only if it contains $z_{i}$; consequently, $M:=\{w_{i}z_{i}|i\in\{0,1,\dots,2k\}{\rm~{}and~{}}\{w_{i},z_{i}\}\cap V(C)=\emptyset\}$ is a perfect matching of $G-V(C)$; thus, $C$ is conformal. This proves the following.
Proposition 26.
Prisms are cycle-extendable.∎
There are two more Norine-Thomas families each of whose member is planar; these are “staircases” and “truncated biwheels”. The smallest member of each of these families is the triangular prism $\overline{C_{6}}$. We refer the reader to Kothari and Murty (2016) for their descriptions — using which they may easily verify the following.
Proposition 27
Every staircase as well as truncated biwheel, except $\overline{C_{6}}$, has a mixed bicycle; consequently, each of them is $K_{2,3}$-based. ∎
To summarize, we have proved the following.
Corollary 28.
[Planar $K_{2,3}$-free Norine-Thomas Bricks]
Wheels and prisms are the only planar Norine-Thomas bricks that are $K_{2,3}$-free; in fact, they are also cycle-extendable.∎
In the next section, we describe the Norine-Thomas brick generation theorem.
4.2 Norine-Thomas brick generation theorem
We first state the induction viewpoint of the Norine-Thomas result using the terminology of Carvalho et al. (2015).
$v_{0}$$v_{1}$$v_{2}$
(a) $G$
$v$
(b) $G/v_{0}$
Figure 8: an illustration of the bicontraction operation
Let $G$ be a matching covered graph, and let $v_{0}$ be a vertex of degree two that has two distinct neighbors $v_{1}$ and $v_{2}$. The bicontraction of $v_{0}$ is the operation of contracting the two edges $v_{0}v_{1}$ and $v_{0}v_{2}$, and we use $G/v_{0}$ to denote the resulting graph; see Figure 8. Note that $\partial(X)$ is a tight cut, where $X:=\{v_{0},v_{1},v_{2}\}$, and $G/v_{0}$ is simply the $\partial(X)$-contraction $G/X$, as defined in Section 2.1; consequently, $G/v_{0}$ is matching covered. It is worth noting that even if $G$ is simple, $G/v_{0}$ need not be simple. The following is easily proved and we shall use its contrapositive in our proof of the characterization of planar $K_{2,3}$-free bricks.
Lemma 29.
Let $G$ be a matching covered graph and let $v_{0}$ denote a vertex of degree two that has two distinct neighbors. For any graph $K$ with maximum degree $\Delta(K)\leq 3$, if $G/v_{0}$ is $K$-based then $G$ is also $K$-based. ∎
Now, let simple $G$ be a brick and $e$ denote a removable edge as defined in the preceding section. Since each vertex of $G$ has three or more distinct neighbors, the matching covered graph $G-e$ is irreducible and it has zero, one or two vertices of degree two. The matching covered graph $J$ obtained from $G-e$ by bicontracting each of its vertices of degree two (if any) is referred to as the retract of $G-e$. Note that its minimum degree $\delta(J)\geq 3$; however, $J$ need not be a brick. For instance, in the Petersen graph $\mathbb{P}$, shown in Figure 9(a), each edge is removable; however, the reader may verify that $b(J)=2$, where $J$ is retract of $\mathbb{P}-e$ for any edge $e$.
$e$
$e$
Figure 9: (a) the Petersen graph $\mathbb{P}$; (b) the bicorn $R_{8}$; (c) the tricorn $R_{10}$.
In light of the above discussion, a removable edge $e$ of a brick $G$ is a thin edge if the retract of $G-e$ is also a brick. For instance, the bicorn $R_{8}$, shown in Figure 9(b), has a unique removable edge $e$ that also happens to be thin — since the retract of $R_{8}-e$ is the brick $K_{4}$ with multiple edges. Carvalho et al. (2006) proved that every brick, except $K_{4}$, triangular prism $\overline{C_{6}}$ and the Petersen graph, has a thin edge.
A thin edge $e$ of a simple brick $G$ is a strictly thin edge if the retract of $G-e$ is also simple. For instance, the tricorn $R_{10}$, shown in Figure 9(c), has precisely three strictly thin edges; for each such edge, say $e$, the retract of $R_{10}-e$ is precisely the wheel $W_{5}$. On the other hand, the bicorn $R_{8}$ has no strictly thin edges. We are now ready to state the generation theorem for simple bricks due to Norine and Thomas (2007).
Theorem 30.
[Strictly Thin Edge Theorem]
Every simple brick, except for the Norine-Thomas bricks, has a strictly thin edge.
In order to prove our characterization of planar $K_{2,3}$-free bricks, we require the generation viewpoint of the above theorem which states that every simple brick may be generated from a Norine-Thomas brick by a sequence of four “expansion operations” as described by Carvalho et al. (2006).
Let $H$ be a graph that has a vertex of degree at least four, say $v$. Consider any graph $G$ that is obtained from $H$ by replacing the vertex $v$ by two new vertices $v_{1}$ and $v_{2}$, distributing the edges of $\partial_{H}(v)$ amongst $v_{1}$ and $v_{2}$ so that each of them receives at least two, and then adding a new vertex $v_{0}$ as well as edges $v_{0}v_{1}$ and $v_{0}v_{2}$. We say that $G$ is obtained from $H$ by bisplitting the vertex $v$. Observe that $H=G/v_{0}$; see Figure 8. It is easily seen that $G$ is matching covered if and only if $H$ is matching covered.
Each of the aforementioned four “expansion operations” consists of bisplitting zero, one or two vertices of a simple brick $J$ and then adding a suitable edge $e$ in order to obtain a larger simple brick $G$. Conversely, $e$ is a strictly thin edge of $G$ and $J$ is the retract of $G-e$. The interested reader may refer to Carvalho et al. (2006) for the exact descriptions of these expansion operations. However, we shall only use the bisplitting operation in our proof of the characterization of planar $K_{2,3}$-free bricks; in particular, we will find the following observation useful.
Corollary 31.
Let $G$ denote a graph obtained from a graph $H$ by bisplitting a vertex of degree four or more. If $H$ contains a mixed bicycle then so does $G$. ∎
The above can also be deduced from Lemma 29 by noting that a mixed bicycle is simply a bisubdivision of the disjoint union of $C_{3}$ and $C_{4}$, and that bicontraction may be viewed as the “inverse” of bisplitting.
4.3 Characterization of planar $K_{2,3}$-free bricks
We are now ready to present our characterization of planar $K_{2,3}$-free bricks.
Theorem 32.
[Planar $K_{2,3}$-free Bricks]
A simple planar brick is $K_{2,3}$-free if and only if it is either a wheel or a prism.
Proof 4.1.
In light of Corollary 28, we only need to prove the forward direction. To this end, let $G$ denote a simple planar brick that is $K_{2,3}$-free. We proceed by induction on the number of edges.
If $G$ is a Norine-Thomas brick then we are done by Corollary 28. Otherwise, by Theorem 30, $G$ has a strictly thin edge, say $e$, and we let $J$ denote the retract of $G-e$. Since $G$ is $K_{2,3}$-free, so is $G-e$; by Lemma 29, so is $J$. Ergo, $J$ is a simple planar brick that is $K_{2,3}$-free; by the induction hypothesis, $J$ is either a wheel or a prism. In particular, $G$ is obtained from $J$ by zero, one or two bisplitting operations followed by adding an edge. We shall consider two cases depending on the number of bisplitting operations — either zero or at least one.
First, suppose that $G$ is obtained from $J$ by adding an edge; that is, $G=J+e$. Since $J$ is either a wheel or a prism, the reader may easily verify that if $|V(J)|\geq 8$ then $G$ contains a mixed bicycle; by Lemma 22, $G$ is $K_{2,3}$-based; contradiction. On the other hand, if $|V(J)|=6$, then $G$ is obtained from $W_{5}$ or $\overline{C_{6}}$ by adding the edge $e$; in each case, one may observe that $K_{2,3}$ is a subgraph of $G$; contradiction.
Now suppose that $G$ is obtained from $J$ by one or two bisplitting operations followed by adding an edge. In particular, $J$ has a vertex of degree four or more. Ergo, $J$ is an odd wheel of order six or more. Furthermore, the first bisplitting operation must bisplit the hub $h$ of $J$, the reader may easily verify that the resulting graph has a mixed bicycle. By Lemma 31, $G$ contains a mixed bicycle. By Lemma 22, $G$ is $K_{2,3}$-based; contradiction. This completes the proof of Theorem 32.
The above theorem, along with Corollary 28, yields the following.
Corollary 33.
[Planar Cycle-Extendable Bricks]
A simple planar brick is cycle-extendable if and only if it is either a wheel or a prism. ∎
We now proceed towards our final goal of characterizing planar cycle-extendable irreducible graphs.
5 Planar cycle-extendable irreducible graphs
In light of Proposition 0.5, Corollary 19 and Theorem 32, it remains to characterize those planar cycle-extendable irreducible graphs that have a vertex of degree two; let $G$ be such a graph and let $x_{0}$ denote any vertex of degree two. In our inductive proof of the Main Theorem (53), we shall consider the smaller graph $J^{{}^{\prime}}:=G/x_{0}$. Observe that $J^{{}^{\prime}}$ is a planar matching covered graph. By Lemma 0.10, $J^{{}^{\prime}}$ is also cycle-extendable. However, $J^{{}^{\prime}}$ need not be irreducible; in fact, $J^{\prime}$ need not be simple. We now introduce the notion of an “osculating bicycle” that will help us in deducing that $J^{{}^{\prime}}$ is either simple, or otherwise has exactly two multiple edges.
5.1 Osculating bicycle
A pair of cycles $(Q,Q^{{}^{\prime}})$ in a graph $G$ is said to be an osculating bicycle if (i) they intersect in precisely one vertex, and (ii) they have the same parities. Furthermore, if each of them is an odd cycle then we refer to $(Q,Q^{{}^{\prime}})$ as an odd osculating bicycle. Likewise, if each of them is an even cycle then $(Q,Q^{{}^{\prime}})$ is an even osculating bicycle.
We recall the following definition from Section 2.3. For a vertex $v$ of a graph $G$, a cycle $C$ (in $G-v$) is $v$-isolating if $v$ is an isolated vertex in the graph $G-V(C)$. The next lemma provides an easy sufficient condition to deduce that a graph (obtained via bisplitting a vertex) is not cycle-extendable. We shall find this lemma immensely useful in our proof of the Main Theorem (53).
Lemma 34.
[Osculating Bicycle Lemma]
Let $G$ be a graph and $x_{0}$ denote a vertex of degree two that has two distinct neighbors. If $J^{{}^{\prime}}:=G/x_{0}$ has an osculating bicycle $(Q,Q^{{}^{\prime}})$ such that neither $Q$ nor $Q^{{}^{\prime}}$ is a cycle in $G$, then $Q\cup Q^{{}^{\prime}}$ is an $x_{0}$-isolating even cycle in $G$.
$x$$Q$$Q^{{}^{\prime}}$
(a) $J^{{}^{\prime}}:=G/x_{0}$
[…]
(b) $G$
Figure 10: an illustration for the Osculating Bicycle Lemma and its proof
Proof 5.1.
We let $x_{1}$ and $x_{2}$ denote the two distinct neighbors of $x_{0}$ in $G$, and let $x$ denote the contraction vertex in $J^{{}^{\prime}}:=G/x_{0}$. Suppose that $(Q,Q^{{}^{\prime}})$ is an osculating bicycle in $J^{{}^{\prime}}:=G/x_{0}$ such that neither $Q$ nor $Q^{{}^{\prime}}$ is a cycle in $G$. Since $Q$ is not a cycle in $G$, the vertex $x$ belongs to $Q$, and one of the two edges of $Q\cap\partial_{J^{{}^{\prime}}}(x)$ is incident with $x_{1}$ in $G$ and the other edge is incident with $x_{2}$ in $G$; see Figure 10. An analogous statement holds for $Q^{{}^{\prime}}$. Since $(Q,Q^{{}^{\prime}})$ is an osculating bicycle in $J^{{}^{\prime}}$, we infer that $Q$ and $Q^{{}^{\prime}}$ meet precisely at the vertex $x$ in $J^{{}^{\prime}}$. This fact, combined with the fact that their parities are the same, implies that $Q\cup Q^{{}^{\prime}}$ is an $x_{0}$-isolating even cycle in $G$, as shown in Figure 10(b).
For distinct vertices $u$ and $v$ of a graph $G$, we let $\mu_{G}(u,v)$ denote the number of edges joining $u$ and $v$. We shall find this notation useful in the proof of the following corollary of the above lemma.
Corollary 35.
Let $G$ be a simple cycle-extendable graph, $x_{0}$ be a vertex of degree two and let $J^{{}^{\prime}}:=G/x_{0}$. Then either $J^{{}^{\prime}}$ is simple or otherwise $J^{{}^{\prime}}$ has precisely two multiple edges. Furthermore, if $G$ is irreducible then the underlying simple graph $J$ of $J^{{}^{\prime}}$ is also irreducible.
Proof 5.2.
Observe that, since $G$ is simple, all multiple edges in $J^{{}^{\prime}}$ (if any) are incident with the contraction vertex that we denote by $x$.
To prove the first part, suppose to the contrary that $J^{{}^{\prime}}$ has more than two multiple edges; we consider two cases. First suppose that there exists a vertex $u\in V(J^{{}^{\prime}})-x$ such that $\mu_{J^{{}^{\prime}}}(x,u)\geq 3$. Since $J^{{}^{\prime}}=G/x_{0}$, it follows from the pigeonhole principle that $G$ is not simple; contradiction. Otherwise, there exist distinct $u_{1},u_{2}\in V(J^{{}^{\prime}})-x$ such that $\mu_{J^{{}^{\prime}}}(x,u_{1})=\mu_{J^{{}^{\prime}}}(x,u_{2})=2$. Let $e$ and $e^{{}^{\prime}}$ denote distinct edges joining $u_{1}$ and $x$; likewise, let $f$ and $f^{{}^{\prime}}$ denote distinct edges joining $u_{2}$ and $x$. Observe that $(Q:=ee^{{}^{\prime}},Q^{{}^{\prime}}:=ff^{{}^{\prime}})$ is an osculating bicycle in $J^{{}^{\prime}}$ and neither $Q$ nor $Q^{{}^{\prime}}$ is a cycle in the simple graph $G$. By the Osculating Bicycle Lemma (34), $Q\cup Q^{{}^{\prime}}$ is an $x_{0}$-isolating $4$-cycle in $G$; this contradicts the hypothesis that $G$ is cycle-extendable. This proves the first part.
Now suppose that $G$ is irreducible and let $x_{1}$ and $x_{2}$ denote the neighbors of $x_{0}$; thus, $d_{G}(x_{1})\geq 3$ and $d_{G}(x_{2})\geq 3$. Consequently, $d_{J^{{}^{\prime}}}(x)\geq 4$ and degree two vertices of $J^{{}^{\prime}}$ comprise a stable set. If $J^{{}^{\prime}}=J$ then we are done. Otherwise, by the first part, there exists a unique vertex $u\in V(J^{{}^{\prime}})-x$ such that $\mu_{J^{{}^{\prime}}}(x,u)=2$; let $e$ and $e^{{}^{\prime}}$ denote the two edges joining $x$ and $u$. Adjust notation so that $J=J^{{}^{\prime}}-e^{{}^{\prime}}$. Since $G$ is simple, adjust notation so that $e\in\partial_{G}(x_{1})$ and $e^{{}^{\prime}}\in\partial_{G}(x_{2})$. Suppose to the contrary that, unlike $J^{{}^{\prime}}$, the vertices of degree two of $J$ do not comprise a stable set. Since $d_{J}(x)\geq 3$, we infer that $d_{J}(u)=2$ and that the neighbour of $u$, distinct from $x$, also has degree two in $J$. We choose $f\in\partial_{G}(x_{1})-e-x_{1}x_{0}$ and $f^{{}^{\prime}}\in\partial_{G}(x_{2})-e^{{}^{\prime}}-x_{2}x_{0}$. Since $f,f^{{}^{\prime}}\in\partial_{J}(x)$, we let $Q$ denote a conformal cycle containing $f$ and $f^{{}^{\prime}}$ in $J$. Since $d_{J}(u)=2$ and $e=ux$, we infer that $u\notin V(Q)$. Thus $(Q,Q^{{}^{\prime}}:=ee^{{}^{\prime}})$ is an osculating bicycle in $J^{{}^{\prime}}$ and neither $Q$ nor $Q^{{}^{\prime}}$ is a cycle in $G$. By the Osculating Bicycle Lemma (34), $Q\cup Q^{{}^{\prime}}$ is an $x_{0}$-isolating cycle in $G$. This contradicts the hypothesis that $G$ is cycle-extendable, and proves the second part.
We now proceed to describe an important class of bipartite cycle-extendable graphs that will play an important role in our description as well as appreciation of planar cycle-extendable irreducible graphs.
5.2 Half biwheels
A half biwheel is any graph $H$ obtained from an even path $P$[$A,B$] with ends, say $u$ and $v$ in $A$, by introducing a new vertex, say $h$, and joining $h$ with each vertex in $A$. Figure 2 shows half biwheels of orders six and eight. The ends of $P$ are called the corners, and $h$ is called the hub, of $H$. Observe that any half biwheel is a cycle-extendable bipartite matching covered graph. Also, we allow $P$ to be $K_{1}$; in this case, $H$ is the smallest half biwheel $K_{2}$, either vertex may be regarded as the hub $h$ and the other one as the (only) corner $u=v$. Note that $C_{4}$ is the second smallest half biwheel; any vertex may be regarded as the hub $h$ and its two distinct neighbors as the corners $u$ and $v$. In all other cases, the hub and (distinct) corners are uniquely determined. The reader may easily verify the following.
Proposition 36.
Every half biwheel $H$ is bipartite and cycle-extendable. Furthermore, each path of $H$, starting at the hub and ending at a corner, is conformal.∎
In each of the following four sections, we describe a family of nonbipartite planar cycle-extendable irreducible graphs $\mathcal{G}_{0},\mathcal{G}_{1},\mathcal{G}_{2}$ and $\mathcal{G}_{3}$. Each member of these families is constructed from $k\geq 1$ half biwheel(s) — by perhaps identifying all of the hubs (only in the case of $\mathcal{G}_{1}$) — and introducing additional vertices (only in the case of $\mathcal{G}_{3}$) as well as edges. For the sake of brevity, we shall find it convenient to use arithmetic modulo $k$. The reader may verify that these families are pairwise disjoint.
Additionally, the reader may verify that each member of these families is a subdivision of a $3$-connected planar graph. Hence, by Whitney’s Theorem (Bondy and Murty, 2008, Theorem 10.28), it admits a unique planar embedding. We will provide planar embeddings of certain small members of these families. Some of our propositions shall refer to the cyclic ordering of edges (incident at a common vertex) induced by the planar embedding.
For each family, we will state propositions pertaining to the existence of specific even cycles and osculating bicycles. We shall find these very useful in our proof of the Main Theorem in order to invoke the Osculating Bicycle Lemma (34) and thus significantly reduce the amount of case analysis required. In particular, propositions pertaining to even cycles shall be useful in the case where $x$ is a cubic vertex of $J$, as per the notation in the proof of Lemma 34 and Figure 10, whereas propositions pertaining to osculating bicycles shall be invoked in the case where $x$ has degree four or more in $J$.
5.3 Generalized prisms $\mathcal{G}_{0}$
In this section, we introduce our first family of graphs, generalized prisms, denoted by $\mathcal{G}_{0}$. Each member of $\mathcal{G}_{0}$, say $J$, is constructed from the union of $k$ disjoint half biwheels, say $H_{i}$[$A_{i},B_{i}$] for $i\in\{0,1,\ldots,k-1\}$, where $k$ is odd and at least three, each labeled as follows — if the hub $h_{i}$ belongs to $A_{i}$ then label $h_{i}$ as $w_{i}=x_{i}$ and the corners as $y_{i}$ and $z_{i}$, and if the hub $h_{i}$ belongs to $B_{i}$ then label $h_{i}$ as $y_{i}=z_{i}$ and the corners as $w_{i}$ and $x_{i}$ — by adding the following edges: $\alpha_{i}:=x_{i}w_{i+1}$ and $\beta_{i}:=y_{i}z_{i+1}$ for each $10\leq i\leq k-$ 1. Figure 11 depicts an example where the half biwheels $H_{i}$ are shown in blue.
$z_{0}$$y_{0}$$w_{0}=x_{0}$$w_{1}$$x_{1}$$y_{1}=z_{1}$$y_{2}=z_{2}$$w_{2}=x_{2}$$w_{3}$$x_{3}$$y_{3}=z_{3}$$y_{4}=z_{4}$$w_{4}=x_{4}$
Figure 11: a generalized prism
Each member $J$ of $\mathcal{G}_{0}$ is a near-bipartite planar (matching covered) graph. Furthermore, $R_{i}:=\{\alpha_{i},\beta_{i}\}$ denotes a removable doubleton for each $10\leq i\leq k-$ 1, and these are the only removable doubletons. The components of $J-R_{0}-R_{1}-\ldots-R_{k-1}$ are precisely the half biwheels $H_{0},H_{1},\ldots,H_{k}$. Observe that if each half biwheel $H_{i}$ is isomorphic to $K_{2}$ then $J$ is simply a prism. In this sense, the family $\mathcal{G}_{0}$ indeed generalizes the prisms. In order to prove that members of $\mathcal{G}_{0}$ are cycle-extendable, we shall find the following lemma useful; the reader may prove it easily.
Lemma 37.
Let $J$ be a graph, let $(V_{0},V_{1},\ldots,V_{k-1})$ denote a partition of $V(J)$, let $H_{i}$ denote the induced subgraph $J[V_{i}]$ for each $10\leq i\leq k-$ 1, and let $L$ denote a subgraph of $J$. If $L_{i}:=L\cap H_{i}$ is a conformal subgraph of $H_{i}$ for each $10\leq i\leq k-$ 1, then $L$ is a conformal subgraph of $J$. ∎
We are now ready to prove that generalized prisms are cycle-extendable.
Proposition 38.
Each member of $\mathcal{G}_{0}$ is cycle-extendable.
Proof 5.3.
Let $J\in\mathcal{G}_{0}$ and let $R_{0},R_{1},\ldots,R_{k-1}$ denote its removable doubletons. Let $H_{0},H_{1},\ldots,H_{k-1}$ denote the (half biwheel) components of $J-R_{0}-R_{1}-\ldots-R_{k-1}$; adjust notation so that $\partial(V(H_{i}))=R_{i-1}\cup R_{i}$ and let $V_{i}:=V(H_{i})$ for each $10\leq i\leq k-$ 1.
Let $C$ denote an even cycle of $J$. Let us pick $i\in\{0,1,\ldots,k-1\}$, and make some observations. By Lemma 25, $|C\cap R_{i}|\in\{0,2\}$. Since $\partial(V(H_{i}))=R_{i-1}\cup R_{i}$, we infer that $C\cap\partial(V(H_{i}))\in\{\emptyset,R_{i-1},R_{i},R_{i-1}\cup R_{i}\}$. Let $C_{i}:=C\cap H_{i}$. Note that $C\cap\partial(V(H_{i}))=\emptyset$ if and only if either $C_{i}=C$ or $C_{i}$ is the null subgraph of $H_{i}$. Secondly, $C\cap\partial(V(H_{i}))\in\{R_{i-1},R_{i}\}$ if and only if $C_{i}$ is a path of the half biwheel $H_{i}$ that starts at the hub and ends at a corner. Lastly, $C\cap\partial(V(H_{i}))=R_{i-1}\cup R_{i}$ if and only if $C_{i}$ is a spanning subgraph of $H_{i}$. By Proposition 36, $C_{i}$ is a conformal subgraph of $H_{i}$. It follows immediately from Lemma 37, with $C$ playing the role of $L$, that $C$ is a conformal cycle of $J$. Thus $J$ is cycle-extendable.
The following property of generalized prisms, pertaining to the existence of even cycles containing a specified cubic vertex but not containing a specified neighbor, is easily verified.
Proposition 39.
[Even Cycles in Generalized Prisms]
Let $J\in\mathcal{G}_{0}$ and let $R_{0},R_{1},\ldots,R_{k-1}$ denote its removable doubletons. Let $x$ denote a cubic vertex and let $H$ denote the (half biwheel) component of $J-R_{0}-R_{1}-\ldots-R_{k-1}$ that contains $x$. Let $w$ denote any neighbor of $x$ in $J$ so that (i) either $w\notin V(H)$ or (ii) $w\in V(H)$ and $d_{J}(w)=2$. Then $J-w$ has an even cycle that contains $x$.
Proof 5.4.
We adopt notation from the definition of generalized prisms, and adjust notation so that $H=H_{1}$. We leave it as an exercise for the reader to locate the desired even cycle in the induced subgraph $J[V(H_{0})\cup V(H_{1})\cup V(H_{2})]$ — by considering various cases depending on whether $H_{1}$ is $K_{2}$ or not, and depending on the specific choices of $x$ and $w$.
For any member $J\in\mathcal{G}_{0}$, it is easy to see that the planar embedding of $J$ has precisely two odd faces whose boundaries are vertex-disjoint and comprise a spanning subgraph of $J$. We use $D$ and $D^{{}^{\prime}}$ to denote these facial cycles. By adjusting notation, one of these cycles contains each $\alpha_{i}$ whereas the other one contains each $\beta_{i}$, where {$R_{i}:=\{\alpha_{i},\beta_{i}\},0\leq i\leq k-1\}$ is the set of all removable doubletons. Using these observation, we now proceed to discuss the existence of specific osculating bicycles in generalized prisms.
$z_{0}$$y_{0}$$w_{0}=x_{0}$$w_{1}=x_{1}$$z_{1}$$t$$t^{{}^{\prime}}$$y_{1}$$f$$f^{{}^{\prime}}$$y_{2}=z_{2}$$w_{2}=x_{2}$$w_{3}$$x_{3}$$y_{3}=z_{3}$$y_{4}=z_{4}$$w_{4}=x_{4}$$\alpha_{0}$$\alpha_{1}$
Figure 12: Illustration for the proof of Proposition 40 (i)
$z_{0}$$y_{0}$$w_{0}=x_{0}$$w_{1}=x_{1}$$z_{1}$$t$$t^{{}^{\prime}}$$y_{1}$$f$$f^{{}^{\prime}}$$y_{2}=z_{2}$$w_{2}=x_{2}$$w_{3}$$x_{3}$$y_{3}=z_{3}$$y_{4}=z_{4}$$w_{4}=x_{4}$$\alpha_{0}$$\alpha_{1}$
Figure 13: Illustration for the proof of Proposition 40 (ii)
Proposition 40.
[Osculating Bicycles in Generalized Prisms]
Let $J\in\mathcal{G}_{0}$ and let $x$ denote a vertex of degree four or more. Let $\alpha_{0}$ and $\alpha_{1}$ denote removable doubleton edges in $\partial(x)$ and let $f,f^{{}^{\prime}}\in\partial(x)-\{\alpha_{0},\alpha_{1}\}$ so that $\alpha_{0},\alpha_{1},f^{{}^{\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Then there exist:
  1. (i)
    an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ such that $\alpha_{0},\alpha_{1}\in E(Q)$ and $f,f^{{}^{\prime}}\in E(Q^{{}^{\prime}})$, and
  2. (ii)
    an even osculating bicycle $(C,C^{{}^{\prime}})$ such that $\alpha_{0},f\in E(C)$ and $\alpha_{1},f^{{}^{\prime}}\in E(C^{{}^{\prime}})$.
Proof 5.5.
We adopt notation from the definition of generalized prisms; thus $x=h_{1}$. Let $f:=h_{1}t$ and $f^{{}^{\prime}}:=h_{1}t^{{}^{\prime}}$. Let $D$ and $D^{{}^{\prime}}$ denote the odd facial cycles as discussed earlier, and adjust notation so that $h_{1}\in V(D)$.
To prove (i), we display an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ whose each constituent cycle uses precisely one edge from each of the removable doubletons of $J$ as follows: $Q:=D$ and $Q^{{}^{\prime}}:=D^{{}^{\prime}}+f+f^{{}^{\prime}}-t(H_{1}-h)t^{{}^{\prime}}$; see Figure 12.
To prove (ii), we display an even osculating bicycle $(C,C^{{}^{\prime}})$ whose each constituent cycle uses either zero or two edges from each of the removable doubletons of $J$ as follows: $C:=h_{1}ft(H_{1}-h)z_{1}\beta_{0}y_{0}x_{0}\alpha_{0}h_{1}$ and $C^{{}^{\prime}}:=h_{1}f^{{}^{\prime}}t^{{}^{\prime}}(H_{1}-h)y_{1}\beta_{1}z_{2}w_{2}\alpha_{1}h_{1}$; see Figure 13.
In each of the following three sections, we will define a family of graphs $\mathcal{G}_{i}$ where $i\in\{1,2,3\}$. Each member of $\mathcal{G}_{i}$ will be defined using some number of half biwheels. For each such half biwheel $H[A,B]$, we adopt the convention that $h$ denotes the hub and that $u$ and $v$ denote the corners (that are not necessarily distinct). Also, this convention is extended naturally to accommodate the use of subscripts.
5.4 Generalized wheels $\mathcal{G}_{1}$
We describe a family of graphs $\mathcal{G}_{1}$. Each member of $\mathcal{G}_{1}$, say $J$, is constructed from $k$ disjoint half biwheels, say $H_{i}$ for $10\leq i\leq k-$ 1, where $k$ is odd and at least three, by identifying all of their hubs into a single vertex $h$ and adding the set of edges $E_{3}:=\{v_{i}u_{i+1}:0\leq i\leq k-1\}$. Figure 14 depicts two examples where the half biwheels $H_{i}$ are shown in blue. We refer to $h$ as a hub of $J$.
Observe that $J-E_{3}$ is precisely the (bipartite) graph obtained by identifying the hubs of $k=|E_{3}|$ half biwheels and $h$ is its unique cut vertex. Furthermore, when $J$ is not $K_{4}$, the set $E_{3}$ comprises precisely those edges whose both ends are cubic. If each half biwheel $H_{i}$ (in the above definition) is isomorphic to $K_{2}$, then $J$ is simply a wheel. In this sense, the family $\mathcal{G}_{1}$ indeed generalizes wheels. Note that the hub $h$ is the unique vertex of degree four or more except when $J$ is $K_{4}$.
$h$$u_{1}=v_{1}$$u_{2}=v_{2}$$v_{0}$$u_{0}$
(a) $W_{5}^{-}$
$u_{0}=v_{0}$$u_{1}$$u_{2}$$v_{1}$$v_{2}$$h$
Figure 14: generalized wheels
Each member $J$ of $\mathcal{G}_{1}$ is a planar matching covered graph and ${\sf f_{odd}}(J)=k+1$. Consequently, they are not near-bipartite in general; see Proposition 16. The graph $J$ is near-bipartite if and only if $k=3$ and at least one of $H_{0}$, $H_{1}$ and $H_{2}$ is isomorphic to $K_{2}$. Furthermore, if $k=3$ and if $H_{i}$ denotes a half biwheel that is isomorphic to $K_{2}$ then $R_{i}:=\{u_{i}h,v_{i+1}u_{i+2}\}$ is a removable doubleton; it follows that the number of half biwheels (among $H_{0}$, $H_{1}$, and $H_{2}$) that are isomorphic to $K_{2}$ equals the number of removable doubletons in $J$.
To see that each member $J$ of $\mathcal{G}_{1}$ is cycle-extendable, note that $J-h$ is an odd cycle; consequently, if $C$ is any even cycle, then $h\in V(C)$ and $J-V(C)$ is an odd path.
Proposition 41.
Each member of $\mathcal{G}_{1}$ is cycle-extendable.∎
We now make an observation pertaining to the existence of even cycles in generalized wheels containing a specified cubic vertex but not containing a specified neighbor.
Proposition 42.
[Even Cycles in Generalized Wheels]
Let $h$ denote the hub of $J\in\mathcal{G}_{1}-K_{4}$ and let $E_{3}$ denote the set of edges whose both ends are cubic. Let $x$ denote a cubic vertex of $J$ and let $w\neq h$ denote a neighbor of $x$ in $J$. If there is no even cycle in $J-w$ that contains $x$, then (i) $|E_{3}|=3$ and (ii) each of $x$ and $w$ is an isolated vertex in $J-E_{3}-h$.
Proof 5.6.
We adopt notation from the definition of generalized wheels, and adjust notation so that $x\in V(H_{1})$ and $w\in V(H_{0})\cup V(H_{1})$. Assume that there is no even cycle in $J-w$ that contains $x$.
First, consider the case in which $w\in V(H_{1})$. If $x\notin\{u_{1},v_{1}\}$, then the reader may locate an even cycle in $H_{1}-w$ that contains $x$; a contradiction. Otherwise, we may adjust notation so that $x=u_{1}$. Observe that $Q:=hu_{1}v_{0}(H_{0}-h)u_{0}v_{k-1}h$ is an even cycle in $J-w$ that contains $x$; a contradiction.
Next, consider the case in which $w\in V(H_{0})$; thus $x=u_{1}$ and $w=v_{0}$. If $H_{1}$ is not $K_{2}$, then $H_{1}$ has an even cycle containing $x$; this contradicts our assumption. Now suppose that $H_{1}$ is $K_{2}$; in other words, $x=u_{1}=v_{1}$. Observe that $Q:=hv_{1}u_{2}(H_{2}-h)v_{2}u_{3}h$ is an even cycle that contains $x$; furthermore, $Q$ does not contain $w=v_{0}$ if and only if $u_{3}\neq v_{0}$. Hence, by our assumption, $u_{3}=v_{0}$. It follows from the definition of generalized wheels that conditions (i) and (ii) hold.
Finally, we proceed to discuss the existence of certain osculating bicycles in generalized wheels.
$h$$u_{0}=v_{0}$$u_{1}$$v_{1}$$u_{2}$$v_{2}$$u_{3}$$v_{3}$$u_{4}=v_{4}$$e_{0}$$e_{1}$$e_{r}$$e_{r+1}$
Figure 15: illustration for the proof of Proposition 43 (iii)
Proposition 43.
[Osculating Bicycles in Generalized Wheels]
Let $h$ denote the hub of $J\in\mathcal{G}_{1}-K_{4}$ and let $E_{3}$ denote the set of edges whose both ends are cubic. Let $H_{0},H_{1},\ldots,H_{k}$ denote the (half biwheel) blocks of $J-E_{3}$ so that they appear in this cyclic order (around the hub $h$) in the planar embedding of $J$, and adjust notation so that $E_{3}=\{v_{i}u_{i+1}:v_{i}\in V(H_{i})~{}{\rm and}~{}u_{i+1}\in V(H_{i+1})\}$. The following statements hold:
  1. (i)
    for any $r$ such that $0\leq r<k$, if the four edges $hu_{0},hv_{r},hu_{r+1}$ and $hv_{k}$ are pairwise distinct then $(Q:=hu_{0}v_{k}h,Q^{{}^{\prime}}:=hv_{r}u_{r+1}h)$ is an odd osculating bicycle in $J$.
  2. (ii)
    for any $0\leq i<r\leq k$, for distinct $e$ and $f$ in $E(H_{i})\cap\partial_{J}(h)$ and distinct $e^{{}^{\prime}}$ and $f^{{}^{\prime}}$ in $E(H_{r})\cap\partial_{J}(h)$, there exists an even osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$ such that $e,f\in E(Q)$ and $e^{{}^{\prime}},f^{{}^{\prime}}\in E(Q^{{}^{\prime}})$.
  3. (iii)
    for any $0<r\leq k$, for any four pairwise distinct edges $e_{0},e_{1},e_{r},e_{r+1}\in\partial_{J}(h)$ that appear in this cyclic order in the planar embedding of $J$, such that $e_{i}\in E(H_{i})$, there exists an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$ such that $e_{0},e_{1}\in E(Q)$ and $e_{r},e_{r+1}\in E(Q^{{}^{\prime}})$.
Proof 5.7.
The reader may easily verify statement (i), whereas statement (ii) follows from the fact that any two adjacent edges (in $H_{i}$; likewise, in $H_{r}$) belong to an even cycle.
We now display an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ for statement (iii). For each $i\in\{0,1,r,r+1\}$, we let $w_{i}$ denote the end of $e_{i}$ that is distinct from $h$. We define them as follows: $Q:=hw_{0}(H_{0}-h)v_{0}u_{1}(H_{1}-h)w_{1}h$ and $Q^{{}^{\prime}}:=hw_{r}(H_{r}-h)v_{r}u_{r+1}(H_{r+1}-h)w_{r+1}h$; see Figure 15. Both cycles use precisely one edge from the set $E_{3}$; thus, they are odd cycles.
5.5 Double half biwheels $\mathcal{G}_{2}$
This section describes another family of graphs denoted as $\mathcal{G}_{2}$. Each member of $\mathcal{G}_{2}$, say $J$, is constructed from the disjoint union of two half biwheels neither of which is isomorphic to $K_{2}$, say $H_{0}$ and $H_{1}$, by adding the following four edges: $\alpha_{0}:=h_{1}h_{0}$, $\beta_{0}:=v_{1}v_{0}$, $\alpha_{1}:=u_{0}h_{1}$, and $\beta_{1}:=h_{0}u_{1}$. Figure 16 depicts an example where the half biwheels $H_{0}$ and $H_{1}$ are shown in blue. (The reader may verify that if any of $H_{0}$ and $H_{1}$ is allowed to be isomorphic to $K_{2}$, then $J$ is a generalized wheel.)
$h_{0}$$v_{0}$$u_{0}$$h_{1}$$u_{1}$$v_{1}$
Figure 16: a double half biwheel
The reader may observe that, in the above definition, the half biwheels $H_{0}$ and $H_{1}$ are interchangeable and that $h_{0}$ and $h_{1}$ are the only vertices of degree four or more. Each member of $\mathcal{G}_{2}$, say $J$ is a near-bipartite planar (matching covered) graph. Furthermore, $R_{0}=\{\alpha_{0},\beta_{0}\}$ and $R_{1}=\{\alpha_{1},\beta_{1}\}$ are the only removable doubletons, and the components of $J-R_{0}-R_{1}$ are precisely the half biwheels $H_{0}$ and $H_{1}$. Using ideas similar to the ones in the proof of Proposition 38, the reader may verify the following.
Proposition 44.
Each member of $\mathcal{G}_{2}$ is cycle-extendable.∎
As usual, we make an observation concerning the existence of even cycles in double half biwheels containing a specified cubic vertex but not containing a specified neighbor. We leave its proof as an exercise for the reader.
Proposition 45.
[Even Cycles in Double Half Biwheels]
Let $J\in\mathcal{G}_{2}$ and let $R_{0},R_{1}$ denote the removable doubletons of $J$. Let $x$ denote a cubic vertex and let $H_{1}$ denote the (half biwheel) component of $J-R_{0}-R_{1}$ that contains $x$. Let $w$ denote a neighbor of $x$ in $J$ so that (i) either $w\notin V(H_{1})$ or (ii) $w\in V(H_{1})$ and $d_{J}(w)=2$. Then, $J-w$ has an even cycle that contains $x$.∎
We now proceed to state two observations pertaining to the existence of osculating bicycles; the first of these deals with vertices of degree four or more.
Proposition 46.
[Osculating Bicycles in Double Half Biwheels - I]
Let $J\in\mathcal{G}_{2}$ and let $x$ denote a vertex of degree four or more. Let $\alpha_{0}$ and $\alpha_{1}$ denote removable doubleton edges in $\partial(x)$ and let $f,f^{{}^{\prime}}\in\partial(x)-\{\alpha_{0},\alpha_{1}\}$ so that $\alpha_{0},\alpha_{1},f^{{}^{\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Then there exists an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ such that $\alpha_{0},f\in E(Q)$ and $\alpha_{1},f^{{}^{\prime}}\in E(Q^{{}^{\prime}})$.
Proof 5.8.
We adopt notation from the definition of double half biwheels; thus $x=h_{1}$. Let $f:=h_{1}w$ and $f^{{}^{\prime}}:=h_{1}w^{{}^{\prime}}$. We display an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ whose each constituent cycle uses precisely one edge from each of the two removable doubletons of $J$ as follows: $Q:=h_{1}fw(H_{1}-h_{1})u_{1}\beta_{1}h_{0}\alpha_{0}h_{1}$ and $Q^{{}^{\prime}}:=h_{1}f^{{}^{\prime}}w^{{}^{\prime}}(H_{1}-h_{1})v_{1}\beta_{0}v_{0}(H_{0}-h_{0})u_{0}\alpha_{1}h_{1}$. Figure 17(a) shows an illustration.
$h_{0}$$v_{0}$$u_{0}$$h_{1}$$u_{1}$$w$$w^{{}^{\prime}}$$v_{1}$$\alpha_{0}$$\alpha_{1}$$f$$f^{{}^{\prime}}$
$h_{0}$$v_{0}$$u_{0}$$h_{1}$$u_{1}=w^{{}^{\prime}}$$w^{{}^{\prime\prime}}$$w$$v_{1}$$\alpha_{0}$$\alpha_{1}$$f^{{}^{\prime}}$$f$$f^{{}^{\prime\prime}}$
Figure 17: illustration for the proofs of Propositions 46 and 47
The next proposition deals with vertices of degree five or more.
Proposition 47.
[Osculating Bicycles in Double Half Biwheels - II]
Let $J\in\mathcal{G}_{2}$ and let $x$ denote a vertex of degree five or more. Let $\alpha_{0}$ and $\alpha_{1}$ denote removable doubleton edges in $\partial(x)$ and let $f,f^{{}^{\prime}},f^{{}^{\prime\prime}}\in\partial(x)-\{\alpha_{0},\alpha_{1}\}$ so that $\alpha_{1},\alpha_{0},f^{{}^{\prime}},f^{{}^{\prime\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Then there exists an even osculating bicycle $(Q,Q^{{}^{\prime}})$ such that $\alpha_{0},f\in E(Q)$ and $f^{{}^{\prime}},f^{{}^{\prime\prime}}\in E(Q^{{}^{\prime}})$.
Proof 5.9.
We adopt notation from the definition of double half biwheels; thus $x=h_{1}$. Let $f:=h_{1}w$, $f^{{}^{\prime}}:=h_{1}w^{{}^{\prime}}$ and $f^{{}^{\prime\prime}}:=h_{1}w^{{}^{\prime\prime}}$. We display an even osculating bicycle $(Q,Q^{{}^{\prime}})$ whose each constituent cycle meets each of the two removable doubletons in zero or two edges as follows: $Q:=h_{1}fw(H_{1}-h_{1})v_{1}\beta_{0}v_{0}h_{0}\alpha_{0}h_{1}$ and $Q^{{}^{\prime}}:=h_{1}f^{{}^{\prime}}w^{{}^{\prime}}(H_{1}-h_{1})w^{{}^{\prime\prime}}f^{{}^{\prime\prime}}h_{1}$. See Figure 17(b) for an example.
5.6 Hexagon half biwheels $\mathcal{G}_{3}$
This section introduces our last family of graphs denoted as $\mathcal{G}_{3}$. A member of $\mathcal{G}_{3}$, say $J$, is obtained from the disjoint union of a hexagon (that is, a $6$-cycle) $H_{0}:=a_{0}a_{1}a_{2}a_{3}a_{4}a_{5}a_{0}$ and a half biwheel $H_{1}$ by adding the following edges: $\alpha_{0}:=h_{1}a_{4}$, $\beta_{0}:=v_{1}a_{1}$, $\alpha_{1}:=a_{3}h_{1}$ and $\beta_{1}:=a_{0}u_{1}$. Figure 18 depicts two examples where $H_{0}$ and $H_{1}$ are shown in blue. Observe that if $H_{1}$ is isomorphic to $K_{2}$ then $J$ is the graph $R_{8}^{-}$.
$a_{0}$$u_{1}$$v_{1}$$a_{1}$$a_{4}$$h_{1}$$a_{3}$$a_{5}$$a_{2}$
$a_{0}$$u=v$$a_{1}$$a_{4}$$h_{1}$$a_{3}$$a_{5}$$a_{2}$
Figure 18: hexagon half biwheels
The reader may easily verify that each member of $\mathcal{G}_{3}$, say $J$, is also near-bipartite (matching covered) graph and $R_{0}=\{\alpha_{0},\beta_{0}\}$, $R_{1}=\{\alpha_{1},\beta_{1}\}$ are the only removable doubletons. Furthermore, the components of $J-R_{0}-R_{1}$ are precisely the $6$-cycle $H_{0}$ and the half biwheel $H_{1}$. Using ideas similar to the ones in the proof of Proposition 38, the reader may verify the following.
Proposition 48.
Each member of $\mathcal{G}_{3}$ is cycle-extendable.∎
We now proceed to make a couple of observations pertaining to the existence of even cycles containing a specified cubic vertex but not containing a specified neighbor; the first of these considers the case in which the cubic vertex belongs to the hexagon. We leave their proofs as exercises for the reader.
Proposition 49.
[Even Cycles in Hexagon Half Biwheels - I]
Let $J\in\mathcal{G}_{3}$ and let $R_{0},R_{1}$ denote the removable doubletons of $J$. Let $H_{0}$ denote the component of $J-R_{0}-R_{1}$ that is isomorphic to a $6$-cycle. Let $x$ denote a cubic vertex such that $x\in V(H_{0})$, and $w$ denote any neighbor of $x$. Then $J-w$ has an even cycle that contains $x$.∎
The next proposition considers the case in which the cubic vertex belongs to the half biwheel.
Proposition 50.
[Even Cycles in Hexagon Half Biwheels - II]
Let $J\in\mathcal{G}_{3}$ and let $R_{0},R_{1}$ denote the removable doubletons of $J$. Let $H_{1}$ denote the component of $J-R_{0}-R_{1}$ that is a half biwheel. Let $x$ denote a cubic vertex such that $x\in V(H_{1})$ and $w$ denote a neighbor of $x$ in $J$ so that (i) either $w\notin V(H_{1})$ or (ii) $w\in V(H_{1})$ and $d_{J}(w)=2$. Then $J-w$ has an even cycle that contains $x$.∎
$a_{0}$$u_{1}$$v_{1}$$a_{1}$$a_{4}$$h_{1}$$a_{3}$$a_{5}$$a_{2}$$\alpha_{0}$$\alpha_{1}$$f$$f^{{}^{\prime}}$
$a_{0}$$u_{1}$$v_{1}$$a_{1}$$a_{4}$$h_{1}$$a_{3}$$a_{5}$$a_{2}$$f$$f^{{}^{\prime}}$$\alpha_{0}$$\alpha_{1}$
Figure 19: illustration for the proof of Proposition 51
Finally, we discuss the existence of specific osculating bicycles in hexagon half biwheels.
Proposition 51.
[Osculating Bicycles in Hexagon Half Biwheels]
Let $J\in\mathcal{G}_{3}$ and let $x$ denote a vertex of degree four or more. Let $\alpha_{0}$ and $\alpha_{1}$ denote removable doubleton edges in $\partial(x)$ and let $f,f^{{}^{\prime}}\in\partial(x)-\{\alpha_{0},\alpha_{1}\}$ so that $\alpha_{0},\alpha_{1},f^{{}^{\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Then there exist:
  1. (i)
    an odd osculating bicycle $(Q,Q^{{}^{\prime}})$ such that $\alpha_{0},\alpha_{1}\in E(Q)$ and $f,f^{{}^{\prime}}\in E(Q^{{}^{\prime}})$, and
  2. (ii)
    an odd osculating bicycle $(C,C^{{}^{\prime}})$ such that $\alpha_{0},f\in E(C)$ and $\alpha_{1},f^{{}^{\prime}}\in E(C^{{}^{\prime}})$.
Proof 5.10.
We adopt notation from the definition of hexagon half biwheels; thus $x=h_{1}$. Let $f:=h_{1}w$ and $f^{{}^{\prime}}:=h_{1}w^{{}^{\prime}}$.
To prove (i), we display an odd osculating bicycle $(Q,Q^{{}^{\prime}})$, whose each constituent cycle uses precisely one edge from each of the two removable doubletons of $J$ as follows: $Q:=h_{1}w(H_{1}-h_{1})u_{1}\beta_{1}a_{0}a_{5}a_{4}\alpha_{0}h_{1}$ and $Q^{{}^{\prime}}:=h_{1}w^{{}^{\prime}}(H_{1}-h_{1})v_{1}\beta_{0}a_{1}a_{2}a_{3}\alpha_{1}h_{1}$. Figure 19(a) shows an illustration.
Likewise, to prove (ii), we display an odd osculating bicycle $(C,C^{{}^{\prime}})$, whose each constituent cycle uses precisely one edge from each of the two removable doubletons of $J$ as follows: $C:=h_{1}\alpha_{1}a_{3}a_{4}\alpha_{0}h_{1}$ and $C^{{}^{\prime}}:=h_{1}w(H_{1}-h_{1})u_{1}\beta_{1}a_{0}a_{1}\beta_{0}v_{1}(H_{1}-h_{1})w^{{}^{\prime}}h_{1}$. Figure 19(b) shows an illustration.
5.7 Main Theorem: planar cycle-extendable  irreducible graphs
The following observation pertaining to all of the families defined in the previous section may be easily verified by the reader; see (Bondy and Murty, 2008, Theorem 10.7, Corollary 10.8).
Proposition 52.
Let $J\in\mathcal{G}_{0}\cup\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$ and let $x$ denote a vertex of degree four or more. Then $J-x$ is a $2$-connected (planar) graph. Consequently, the neighbors of $x$ lie on a cycle of $J-x$. ∎
We are now ready to state and prove the Main Theorem.
Theorem 53.
[Planar Cycle-Extendable Irreducible Graphs]
A planar irreducible matching covered graph $G$ is cycle-extendable if and only if $G$ belongs to $\{K_{2}\}\cup\mathcal{G}_{0}\cup\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$.
Proof 5.11.
Propositions 38, 41, 44 and 48 prove the reverse implication. For the forward implication, let $G$ be a planar irreducible matching covered graph that is cycle-extendable. We proceed by induction on the number of edges. Firstly, if $\delta(G)\geq 3$ then, by Corollary 19 and Corollary 33, $G$ is either a wheel or a prism; thus $G\in\mathcal{G}_{0}\cup\mathcal{G}_{1}$.
Now suppose that $G$ has a vertex of degree two, say $x_{0}$, and let $x_{1}$ and $x_{2}$ denote its neighbors. Let $J^{{}^{\prime}}:=G/x_{0}$ and let $x$ denote its bicontraction vertex. Since $G$ is irreducible, $d_{J^{{}^{\prime}}}(x)\geq 4$. Since $G$ is cycle-extendable, by the first part of Corollary 35, $J^{{}^{\prime}}$ has at most two multiple edges. Clearly, multiple edges (if any) are incident at $x$. In case $J^{{}^{\prime}}$ is not simple we use $e$ and $e^{{}^{\prime}}$ to denote its multiple edges. We let $J$ denote the underlying simple graph of $J^{{}^{\prime}}$. Furthermore, either $J=J^{{}^{\prime}}$ or otherwise we adjust notation so that $J=J^{{}^{\prime}}-e^{{}^{\prime}}$. Thus, $d_{J}(x)\geq 3$. Since $G$ is irreducible, by the second part of Corollary 35, $J$ is also irreducible.
Clearly, $J$ and $J^{{}^{\prime}}$ are planar matching covered graphs. By Lemma 0.10, they are also cycle-extendable. Since $J$ is irreducible and $d_{J}(x)\geq 3$, by the induction hypothesis, ${J}$ belongs to $\mathcal{G}_{0}\cup\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$.
We will divide the proof into cases depending on the degree of vertex $x$ and whether $J$ belongs to $\mathcal{G}_{0}$, $\mathcal{G}_{1}$, $\mathcal{G}_{2}$ or $\mathcal{G}_{3}$. Within each case, we will consider various subcases. In each subcase, we will either display an osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J^{{}^{\prime}}$ such that neither of its constituent cycles is a cycle in $G$ and invoke Lemma 34 to arrive at a contradiction (by inferring that $G$ is not cycle-extendable), or otherwise conclude that $G$ belongs to $\mathcal{G}_{0},\mathcal{G}_{1},\mathcal{G}_{2}$ or $\mathcal{G}_{3}$. Depending on the family that $J$ belongs to, we shall adopt the following notation.
Notation 54
If $J\in\mathcal{G}_{0}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$, then we let $R_{0},R_{1},\ldots,R_{k}$ denote all of the removable doubletons of $J$, and $H_{0},H_{1},\ldots,H_{k}$ denote the components of $G-R_{0}-R_{1}-\ldots-R_{k}$. Furthermore, if $J\in\mathcal{G}_{3}$, then we let $H_{0}$ denote the component that is a $6$-cycle.
Notation 55
If $J\in\mathcal{G}_{1}-K_{4}$, then we let $E_{3}$ denote the set of edges whose both ends are cubic, $h$ denote the unique cut vertex of $J-E_{3}$, and $H_{0},H_{1},\ldots,H_{k}$ denote the (half biwheel) blocks of $J-E_{3}$ so that they appear in this cyclic order (around the hub $h$) in the planar embedding of $J$ and adjust notation so that $E_{3}=\{v_{i}u_{i+1}:v_{i}\in V(H_{i})~{}{\rm and}~{}u_{i+1}\in V(H_{i+1})\}$.
Case 1: $d_{J}(x)=3$.
Observe that $J^{{}^{\prime}}$ is obtained from $J$ by adding a multiple edge $e^{{}^{\prime}}$ incident with $x$. Let $e=e^{{}^{\prime}}=xw$ in $J^{{}^{\prime}}$ and let $Q^{{}^{\prime}}$ denote $2$-cycle $ee^{{}^{\prime}}$. Observe that, since $d_{J^{{}^{\prime}}}(x)=4$ and since $G$ is simple, there is a unique way to bisplit $x$ and obtain the graph $G$ from $J^{{}^{\prime}}$. Consequently, if there exists an even cycle $Q$ in $J^{{}^{\prime}}$ such that $(Q,Q^{{}^{\prime}})$ is an osculating bicycle in $J^{{}^{\prime}}$, then neither $Q$ nor $Q^{{}^{\prime}}$ is a cycle in $G$; in this case, by Lemma 5.1, we arrive at a contradiction to the assumption that $G$ is cycle-extendable. Observe that $Q$ exists if and only if $J-w$ has an even cycle that contains $x$. From the preceding discussion, it suffices to consider only those cases in which $J-w$ has no even cycle that contains $x$. In the following two paragraphs, we shall heavily exploit this observation.
If $J\in\mathcal{G}_{l}$ where $l\in\{0,2,3\}$, we adopt Notation 54. If $l=3$, we invoke Proposition 49 to infer that $x\notin V(H_{0})$. Now, we may adjust notation so that $x\in V(H_{1})$. Depending on the value of $l$, we invoke Proposition 39, or 45, or 50, to infer that $d_{J}(w)\geq 3$ and that $w$ is a hub of the same half biwheel $H_{1}$. Since the bisplitting is unique (as noted earlier), the reader may easily verify that $G$ belongs to the same family $\mathcal{G}_{l}$.
Now consider the case in which $J\in\mathcal{G}_{1}$. If $J=K_{4}$ then the reader may easily verify that $G=W_{5}^{-}$. Now suppose that $J\neq K_{4}$ and adopt Notation 55. We invoke Proposition 42 to infer that either $w=h$, or otherwise, $|E_{3}|=3$ and $J-E_{3}-h$ has precisely two isolated vertices — namely, $x$ and $w$. If $w=h$, then using the fact that bisplitting of $x$ is unique, one may easily see that $G\in\mathcal{G}_{1}$. We consider the remaining case below.
$x$$w$$h$
(a) J
$x_{1}$$x_{0}$$x_{2}$$w$$h$
(b) G
Figure 20: the case in which $J\in\mathcal{G}_{1}$ and $G\in\mathcal{G}_{2}$
Suppose that $|E_{3}|=3$ and that $J-E_{3}-h$ has precisely two isolated vertices — namely, $x$ and $w$. Based on the description of $\mathcal{G}_{1}$, this is equivalent to saying that $k=2$ and precisely two of the half biwheels $H_{0}$, $H_{1}$ and $H_{2}$ are isomorphic to $K_{2}$. We adjust notation so that each of $H_{1}$ and $H_{2}$ is isomorphic to $K_{2}$ and that $x\in V(H_{1})$ and $w\in V(H_{2})$. Since the bisplitting is unique, observe that $H:=G$[$w,x_{1},x_{0},x_{2}$] is isomorphic to $C_{4}$, and that $G$ is a member of $\mathcal{G}_{2}$ — wherein $H$ (with hub $w$) and $H_{0}$ (with hub $h$) are the half biwheels as per the definition of double half biwheels in Section 5.5. See Figure 20.
Case 2: $d_{J}(x)\geq 4$.
Note that, unlike the previous case, we must consider all possible bisplittings of the vertex $x$ that lead to a planar graph (instead of just one bisplitting). Observe that $\partial_{J^{{}^{\prime}}}(x)=\partial_{G}(\{x_{0},x_{1},x_{2}\})$. It follows from Proposition 52 that the cyclic order of these edges must be the same in $G$ around the set $\{x_{0},x_{1},x_{2}\}$ as the cyclic order in $J^{{}^{\prime}}$; otherwise, it is easy to see that the resulting graph has a subdivision of $K_{3,3}$.
We consider subcases depending on whether $J$ belongs to $\mathcal{G}_{0}$cup $\mathcal{G}_{2}$cup $\mathcal{G}_{3}$ or to $\mathcal{G}_{1}$. Note that an osculating bicycle in $J$ also exists in $J^{{}^{\prime}}$. It is for this reason that, when applicable, we simply display an osculating bicycle in $J$ with the desired properties in order to arrive at a contradiction as discussed earlier.
Case 2.1: $J\in\mathcal{G}_{0}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$.
We adopt Notation 54. Since $d_{J}(x)\geq 4$, we may adjust notation so that $x$ is a hub of $H_{1}$ that is not isomorphic to $K_{2}$. Let $\alpha_{0}$ and $\alpha_{1}$ denote the removable doubleton edges in $\partial_{J}(x)$. Since $G$ is obtained from $J^{{}^{\prime}}$ by bisplitting the vertex $x$, the edges $\alpha_{0}$ and $\alpha_{1}$ may or may not be adjacent in $G$; we consider these cases separately.
First suppose that $\alpha_{0}$ and $\alpha_{1}$ are adjacent in $G$ and adjust notation so that $\alpha_{0},\alpha_{1}\in\partial_{G}(x_{1})$. We choose $f,f^{{}^{\prime}}\in\partial_{G}(x_{2})-x_{2}x_{0}$ so that $\alpha_{0},\alpha_{1},f^{{}^{\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Depending on whether $J$ belongs to $\mathcal{G}_{0},\mathcal{G}_{2}$ or $\mathcal{G}_{3}$, we invoke Proposition 40 (ii), 46 or 51 (ii), in order to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$.
Now suppose that $\alpha_{0}$ and $\alpha_{1}$ are nonadjacent in $G$ and adjust notation so that $\alpha_{0}\in\partial(x_{1})$ and $\alpha_{1}\in\partial(x_{2})$. We consider two subcases depending on whether $J\in\mathcal{G}_{0}\cup\mathcal{G}_{3}$ or $J\in\mathcal{G}_{2}$. If $J\in\mathcal{G}_{0}\cup\mathcal{G}_{3}$, we choose $f\in\partial_{G}(x_{1})-x_{1}x_{0}-\alpha_{0}$ and $f^{{}^{\prime}}\in\partial_{G}(x_{2})-x_{2}x_{0}-\alpha_{1}$ so that $\alpha_{0},\alpha_{1},f^{{}^{\prime}},f$ appear in this cyclic order in the planar embedding of $J$. Depending on whether $J$ belongs to $\mathcal{G}_{0}$ or $\mathcal{G}_{3}$, we invoke Proposition 40 (i) or 51 (i), in order to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$. Henceforth $J\in\mathcal{G}_{2}$; we consider two subcases depending on whether $d_{J}(x)=4$ or $d_{J}(x)\geq 5$.
Let us first suppose that $d_{J}(x)=4$. Equivalently, $H_{1}$ is isomorphic to a $4$-cycle, say $h_{1}u_{1}yv_{1}h_{1}$, where $h_{1}:=x$. Since $\alpha_{0}\in\partial_{G}(x_{1})$ and $\alpha_{1}\in\partial_{G}(x_{2})$, observe that $x_{1}u_{1}yv_{1}x_{2}x_{0}x_{1}$ is a $6$-cycle in $G$ and that $G\in\mathcal{G}_{3}$ as per the definition of hexagon half biwheels in Section 5.6; see Figure 21.
$h_{0}$$v_{0}$$u_{0}$$h_{1}=x$$u_{1}$$y$$v_{1}$
(a) J
$v_{1}$$v_{0}$$u_{0}$$x_{2}$$u_{1}$$h_{0}$$x_{1}$$y$$x_{0}$
(b) G
Figure 21: the case in which $J\in\mathcal{G}_{2}$ and $G\in\mathcal{G}_{3}$
Lastly, suppose that $d_{J}(x)\geq 5$; consequently, at least one of $x_{1}$ and $x_{2}$ has degree four or more. Adjust notation so that $d_{G}(x_{2})\geq 4$. Now, we may choose $f^{{}^{\prime}}\in\partial_{G}(x_{1})-x_{1}x_{0}-\alpha_{0}$ and distinct $f^{{}^{\prime\prime}},f\in\partial_{G}(x_{2})-x_{2}x_{0}-\alpha_{1}$ so that $\alpha_{1},\alpha_{0},f^{{}^{\prime}},f^{{}^{\prime\prime}},f$ appear in this cyclic order in the planar embedding of $J$. We invoke Proposition 47 to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$.
Case 2.2: $J\in\mathcal{G}_{1}.$
We adopt Notation 55. Since $d_{J}(x)\geq 4$, the vertex $x$ is precisely the hub $h$ of $J$. Note that, for each $0\leq i\leq k$, the edge set of the half biwheel $H_{i}$ (in $J$) may or may not form a half biwheel in $G$. Furthermore, $E(H_{i})$ forms a half biwheel in $G$ if and only if all edges in $E(H_{i})\cap\partial_{J}(x)$ are incident with (precisely) one of $x_{1}$ and $x_{2}$ (in G); in this case, we say that the half biwheel $H_{i}$ remains intact; otherwise we say that the half biwheel $H_{i}$ is destroyed. It follows from planarity that the number of half biwheels that are destroyed is either zero, two or one; we shall consider these cases separately in that order.
First we consider the case in which zero half biwheels are destroyed. We may adjust notation so that there exists $0\leq r<k$ such that: (i) for each $0\leq i\leq r$, all edges in $E(H_{i})\cap\partial_{J}(x)$ are incident with $x_{1}$ (in $G$), and likewise (ii) for each $r<i\leq k$, all edges in $E(H_{i})\cap\partial_{J}(x)$ are incident with $x_{2}$. It follows from the irreducibility of $G$ that the four edges $xu_{0},xv_{r},xu_{r+1}$ and $xv_{k}$ are pairwise distinct. We invoke Proposition 43 (i) to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$.
Next we consider the case in which two half biwheels, say $H_{i}$ and $H_{r}$, are destroyed. Let $e$ denote an edge in $E(H_{i})\cap\partial_{J}(x)$ that is incident with $x_{1}$ (in $G$) and $f$ denote an edge in $E(H_{i})\cap\partial_{J}(x)$ that is incident with $x_{2}$. Likewise, let $e^{{}^{\prime}}$ denote an edge in $E(H_{r})\cap\partial_{J}(x)$ that is incident with $x_{1}$ and $f^{{}^{\prime}}$ denote an edge in $E(H_{r})\cap\partial_{J}(x)$ that is incident with $x_{2}$. We invoke Proposition 43 (ii) to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$.
Lastly we consider the case in which precisely one half biwheel, say $H_{0}$, is destroyed. we may adjust notation so that there exists $0<r\leq k$ such that: (i) for each $0<i\leq r$, all edges in $E(H_{i})\cap\partial_{J}(x)$ are incident with $x_{1}$ (in $G$), and likewise (ii) for each $r<i\leq k$, all edges in $E(H_{i})\cap\partial_{J}(x)$ are incident with $x_{2}$. Unless $H_{0}$ is isomorphic to $C_{4}$, $k=2$, each of $H_{1}$ and $H_{2}$ is isomorphic to $K_{2}$, and $r=1$, the reader may verify that it is possible to choose $e_{0},e_{1},e_{r},e_{r+1}\in\partial_{J}(h)$ that satisfy the conditions stated in Proposition 43 (iii) so that $e_{0}$ and $e_{1}$ are nonadjacent in $G$ and, likewise, $e_{r}$ and $e_{r+1}$ are nonadjacent in $G$; we invoke Proposition 43 (iii) to locate the desired osculating bicycle $(Q,Q^{{}^{\prime}})$ in $J$.
Now suppose that $H_{0}$ is isomorphic to $C_{4}$, $k=2$, each of $H_{1}$ and $H_{2}$ is isomorphic to $K_{2}$, and $r=1$. Observe that $J$ is the graph $W_{5}^{-}$ as per the labels shown in Figure 14(a). First suppose that $J\neq J^{{}^{\prime}}$; up to symmetry, either $e^{{}^{\prime}}=hu_{1}$ or otherwise $e^{{}^{\prime}}=hu_{0}$. In the former case, since $H_{0}$ is destroyed and $G$ is simple, $(Q:=H_{0},Q^{{}^{\prime}}:=ee^{{}^{\prime}})$ is the desired even osculating bicycle in $J^{{}^{\prime}}$. In the latter case, it follows from irreducibility and planarity that $(Q:=hv_{0}u_{2}u_{1}h,Q^{{}^{\prime}}:=ee^{{}^{\prime}})$ is the desired even osculating bicycle in $J^{{}^{\prime}}$. Finally, if $J=J^{{}^{\prime}}$, then it follows from the facts that $H_{0}$ is destroyed and $G$ is irreducible, that $G$ is the graph $R_{8}^{-}\in\mathcal{G}_{3}$ shown in Figure 18(b).
This completes our proof of the Main Theorem (53).
We conclude our paper with the following corollaries of our Main Theorem (53) and Proposition 0.5.
Corollary 56.
A planar bipartite matching covered graph $G$ is cycle-extendable if and only if any irreducible graph H, obtained from $G$ by repeated applications of series and parallel reductions, is isomorphic to $K_{2}$.∎
Corollary 57.
A planar nonbipartite matching covered graph $G$ is cycle-extendable if and only if any irreducible graph H, obtained from $G$ by repeated applications of series and parallel reductions, belongs to $\mathcal{G}_{0}\cup\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$.∎